### v2

# Job-SDF: A Multi-Granularity Dataset for Job Skill Demand Forecasting and Benchmarking

Abstract
In a rapidly evolving job market, skill demand forecasting is crucial as it enables policymakers and businesses to anticipate and adapt to changes, ensuring that workforce skills align with market needs, thereby enhancing productivity and competitiveness. Additionally, by identifying emerging skill requirements, it directs individuals towards relevant training and education opportunities, promoting continuous self-learning and development. However, the absence of comprehensive datasets presents a significant challenge, impeding research and the advancement of this field. To bridge this gap, we present Job-SDF, a dataset designed to train and benchmark job-skill demand forecasting models. Based on 10.35 million public job advertisements collected from major online recruitment platforms in China between 2021 and 2023, this dataset encompasses monthly recruitment demand for 2,324 types of skills across 521 companies. Our dataset uniquely enables evaluating skill demand forecasting models at various granularities, including occupation, company, and regional levels. We benchmark a range of models on this dataset, evaluating their performance in standard scenarios, in predictions focused on lower value ranges, and in the presence of structural breaks, providing new insights for further research. Our code and dataset are publicly accessible via the https://github.com/Job-SDF/benchmark.
1 Introduction
Job skills encompass a range of abilities and competencies essential for performing tasks effectively in the workplace. These skills are broadly categorized into hard skills, such as technical and analytical abilities, and soft skills, including communication, teamwork, and adaptability [1]. Accurate forecasting of skill demand helps businesses and policymakers anticipate and address skill shortages and mismatches, and promotes skill development in high-demand areas, thereby supporting economic growth and stability [2]. By identifying emerging skill requirements, individuals are directed towards relevant training and education opportunities, fostering continuous self-learning and development to stay competitive in the labor market [3]. Moreover, forecasting informs educational and training programs, ensuring that curricula align with the labor market’s evolving needs [4].
Traditionally, skill demand analysis has relied on labor-intensive, survey-based methods limited to specific companies or occupations [5, 6, 7]. However, over the past decade, the rapid evolution of the internet has spurred the emergence of online recruitment platforms. These platforms have become the primary channels for job advertisements for numerous enterprises and organizations, accumulating vast amounts of job advertisement data. By leveraging this data, researchers have formulated skill demand forecasting as a time series task, utilizing various machine learning models such as autoregressive integrated moving average (ARIMA) [8], recurrent neural networks (RNNs) [9], and dynamic graph autoencoders (DyGAEs) [10], to predict future skill needs.
A major challenge impeding progress in this field is the lack of comprehensive and publicly accessible datasets. Existing studies do not provide open-source datasets, making it difficult for researchers to replicate experimental results and identify bottlenecks in current research. Furthermore, these datasets primarily focus on predicting skill demand variations across different occupations, with a notable lack of modeling and prediction at other granularities, such as companies or regions. This limitation hinders comprehensive comparisons between different models and impedes the exploration of potential downstream applications, such as human capital strategy development and regional policy formulation. Additionally, the significant variations in skill demand present further challenges. Existing studies, which rely on metrics such as Mean Squared Error (MSE), struggle to evaluate the performance of skill demand forecasting models for low-frequency skill terms. For instance, some emerging skills, such as large language models (LLMs), may initially show low demand but are crucial for the job market due to their potential to reshape existing occupations.
To this end, in this paper, we propose Job-SDF, a multi-granularity dataset designed for job skill demand forecasting research. Specifically, we collected 10.35 million public job advertisements from major online recruitment platforms in China, posted between January 2021 and December 2023. By extracting skill terms from job advertisement texts, we quantified the monthly skill demand at various granularities, including occupations, companies, and regions, to construct our dataset. This dataset encompasses 2,324 types of skills, 52 occupations, 521 companies, and 7 regions. We then use the Job-SDF dataset to benchmark a wide range of models for job skill demand forecasting tasks at various granularities. These models include statistical time series models (e.g., ARIMA [11]), deep learning-based methods such as RNN-based models [12, 10], Transformer-based models [13, 14, 15, 16, 17], MLP-based models [18, 19], as well as several state-of-the-art time-series forecasters [20, 21]. Performance is evaluated using regression metrics such as Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE). Additionally, we use the Symmetric Mean Absolute Percentage Error (SMAPE) [22] and Relative Root Mean Squared Error (RRMSE) [23] metrics to account for the significantly varying nature of skill demand values, which is particularly useful for evaluating model performance in predicting lower value ranges. Moreover, we further investigate the impact of structural breaks in job skill demand time series data on model performance. The Job-SDF dataset, along with data loaders, example codes for different models, and evaluation setup, are publicly available in our GitHub repository: https://github.com/Job-SDF/benchmark.
2 Related Work
Skill demand forecasting can analyze how skills evolve over time, aiding experts in evaluating technological advancements [24, 25], assessing wage inequality [26, 27], and generating employment opportunities [28]. Furthermore, the skills required in the 21st-century workplace will differ significantly from those in previous eras [29]. Predicting skill demands benefits personal career transitions and corporate management strategies.
Recently, with the rapid accumulation of data and continuous advancements in technology, skill demand forecasting has demonstrated significant vitality. Das et al. proposed a method for dynamic task allocation to investigate the evolution of job task requirements over a decade of AI innovation across different salary levels  [8]. Given the effectiveness of RNN in multi-step prediction, some researchers have integrated skill demand forecasting with RNN algorithms, achieving promising results  [9, 12]. In addition, considering the supply-demand dynamics of the labor market concurrently, CHGH designed a joint prediction model based on the encoder-decoder architecture to achieve trend prediction for both skill supply and demand sides [10]. Moreover, to capture the dynamic information of occupations, a pretraining-enhanced dynamic graph autoencoder has been developed to efficiently forecast skill demand at the occupation granularity [30].
However, the predominance of closed-source datasets has significantly elevated the barrier of researchers and constrained the pace of methodological advancements. Additionally, the current research data focuses either on macro-market skill demand predictions or on analyses at a specific granularity, neglecting multi-level labor market analysis. This limitation generally hampers the transferability of the modeling approaches.
3 Job-SDF Dataset
The Job-SDF dataset is built from job advertisements collected on online recruitment platforms, encompassing dynamic job skill demand time series data at various granularities, recorded monthly from 2021 to 2023. The dataset is CC BY-NC-SA 4.0 licensed, accessible via the URL https://github.com/Job-SDF/benchmark. We summarize the dataset construction process, task description, and dataset analysis below, with supplementary details provided in the Appendix.
3.1 Data Collection and Processing
Job Advertisement Collection. We collected public job advertisements for 52 occupations from 521 companies  [^1] Representative companies were selected based on the number of job postings they published in early 2021. on major online recruitment platforms in China, spanning from January 2021 to December 2023. After removing identical job advertisements posted simultaneously by different companies on various platforms, we obtained 10.35 million unique records. Each record contains five types of information: (1) Job Requirement, which is a text segment that outlines the specific skills required of candidates applying for the job; (2) Company, which identifies the company that posted the job advertisement; (3) Occupation, which specifies the job advertisement’s category. Our dataset encompasses 52 detailed occupations (L2-level), such as front-end development engineer and financial investment analyst. Additionally, these 52 occupations are grouped into 14 broader categories (L1-level); (4) Region, which indicates the primary geographic divisions in China where the job postings are located. These regions are classified based on their geographical orientation and include Central China, South China, North China, East China, Northeast China, Southwest China, and Northwest China; (5) Posting Time, which records the date when the job was posted, including the year, month, and day.
Job Skill Extraction. After acquiring the job advertisement data, we utilized a Named Entity Recognition (NER) model, as referenced in [31], to explicitly extract skill requirements from the Job Requirement of each advertisement.
Specifically, we first annotated a dataset for training the NER model by identifying skill terms within the job requirement texts. To achieve this, we devised a set of regular expressions tailored to the characteristics of skill descriptions and used these to match skill words in job advertisements. Subsequently, we merged all matched skill words to formulate a raw skill dictionary, including their corresponding frequencies across job advertisements. We then filtered out low-frequency words and manually annotated the raw skill dictionary to create a refined skill dictionary. Along this line, we excluded unreasonable skill words matched by the regular expressions that did not appear in the refined skill dictionary, establishing an initial correspondence between the Job Requirement and the skill requirements.
Based on this annotated data, we trained an NER model to extract required skills from the Job Requirement section for all job advertisements. Experts then aggregated the skills extracted by the NER model based on their meaning and content, grouping together those with similar meanings or repeated expressions. This process resulted in a skill dictionary $\mathcal{S}$ of 2,324 standardized skill words, mapping original skill word descriptions to standardized skill words. The skill dictionary was then used to filter and map the skill words extracted by the NER model, ultimately obtaining standardized skill requirements for each job requirement. These standardized requirements were added to the job advertisement data as a new field, Skill Requirements.
Job Skill Demand Estimation. Generally, the demand for different skills in the job market can be estimated by the volume of job advertisements listing these specific skills as requirements within a given time period [10]. Formally, given job advertisement data $\mathcal{P}=\{\mathcal{P}_{1},...,\mathcal{P}_{t},...,\mathcal{P}_{T}\}$, where each $\mathcal{P}_{t}$ represents the job advertisements posted at timestamp $t$, we use $D_{s,t}=\sum_{p\in\mathcal{P}_{t}}\mathbf{1}(s\in p)$ to estimate the demand for skill $s\in\mathcal{S}$ at time $t$. $s\in p$ indicates that job advertisement $p$ requires skill $s$.
Along this line, we can calculate skill demand at various granularities, such as occupation and company levels. We define the sets of L1-level occupations, L2-level occupations, companies, and regions as $\mathcal{A}^{o_{1}}$, $\mathcal{A}^{o_{2}}$, $\mathcal{A}^{c}$, and $\mathcal{A}^{r}$, respectively. The demands for skill $s$ at time $t$ under granularity $i\in\{o_{1},o_{2},c,r\}$ is then defined as follows:
$D_{s,t}^{i}=[D_{s,t,a^{i}}^{i}]_{a^{i}\in\mathcal{A}^{i}},\ \ D_{s,t,a^{i}}^{i}=\sum_{p\in\mathcal{P}_{t}}\mathbf{1}(s\in p)\cdot\mathbf{1}(a^{i}\in p),$ (1)
where $a^{i}\in p$ represents a job advertisement $p$ containing the attribute $a^{i}$ under granularity $i$. Similarly, we can further define skill demands $D_{s,t}^{i,j,...,k}$ across multiple granularities $\{i,j,...,k\}$ by calculating:
$D_{s,t,\overline{a}}^{i,j,...,k}=\sum_{p\in\mathcal{P}_{t}}\mathbf{1}(s\in p)\cdot\mathbf{1}(a^{i}\in p\land a^{j}\in p\land...\land a^{k}\in p),$ (2)
where $\overline{a}=\{a^{i},a^{j},...,a^{k}\}$, $a^{i}\in\mathcal{A}^{i},a^{j}\in\mathcal{A}^{j},...,a^{k}\in\mathcal{A}^{k}$, and $D^{i,j,\ldots,k}_{s,t}\in\mathbb{R}^{|\mathcal{A}^{i}||\mathcal{A}^{j}|\ldots|\mathcal{A}^{k}|}$.
3.2 Job Skill Demand Forecasting Tasks
We study model performance through job skill demand forecasting tasks at different granularities, including single and multiple levels. The primary goal of these tasks is to predict future job skill demands based on historical time series data of various skills. Formally, we have:
Definition 1 (Job Skill Demand Forecasting)
Given a granularity or a set of granularities $g$ and the observed job skill demand series from the previous $K$ timestamps, i.e., $\{D^{g}_{:,t-K+1},...,D^{g}_{:,t}\}$, the goal of job skill demand forecasting is to learn a forecasting model $\mathcal{M}$ to predict the demand values for the next $H$ timestamps, denoted by $\{\hat{D}^{g}_{:,t+1},\ldots,\hat{D}^{g}_{:,t+H}\}$.
Our dataset includes skill demand time series data for L1-level occupations, L2-level occupations, companies, regions, and their combinations. The data spans from 2021 to 2023, with each time slice representing one month. We follow a standard protocol [32] that categorizes all time-series data into training, validation, and test sets in chronological order with a ratio of 9:1:2. In the main text, we demonstrate results with $K$ set to 6 months and consider $H$ as 3 months to evaluate the performance of different forecasting models. More settings and results can be found in the appendix and our project repository. Based on the Job-SDF dataset, other researchers can easily adjust the parameters to suit their research objectives.
(a) The skill demands under two occupations.
(b) Appying the Chow test to two skill demand time series.
Figure 1: Data analysis on Job-SDF. (a) illustrates the long-tail phenomenon of skill demands under the product manager and doctor occupations. (b) illustrates the results under the Chow test for the absence (left) and presence (right) of structural breaks.
3.3 Dataset Analysis
Varying Nature of Skill Demand. The values of skill demand exhibit significant differences and generally follow a long-tail distribution. This indicates that, at a specific granularity, only a few skills have high demand, while a wide range of skills are required by a limited number of jobs. For instance, Figure 1(a) presents the skill demands under the product manager and doctor occupations in December 2022. The results clearly demonstrate the varying nature of skill demand values. This suggests that relying solely on metrics like RMSE to evaluate forecasting models’ performance may overlook the prediction accuracy for low-frequency skills.
Structural Break Phenomenon. As the labor market evolves, job skills that are not widely required today may become crucial in the future, while those currently in high demand may be supplanted by others. This dynamic can induce significant changes in the statistical properties of skill demand time series at various points in time. These changes may be reflected in the mean, variance, trend, or autocorrelation structure of the series. This phenomenon is known as structural breaks. A common method for detecting structural breaks is the Chow test, which evaluates whether there are significant differences in the regression coefficients across different periods [33]. Figure 1(b) illustrates the application of the Chow test in detecting structural breaks in various skill demand time series. The presence of structural breaks can impact the predictive accuracy of forecasting models. Further discussion will be provided in the experimental section.
Figure 2: Pearson Correlation Coefficients.
Inter-Series Correlation. Intuitively, the proposed job skill demand forecasting tasks can be categorized as multivariate time-series forecasting tasks [34]. Figure 2 shows the absolute values of the Pearson correlation coefficients of different skill demand series for the backend development engineer, salesperson, and product manager occupations. We found that the time series data for some skills exhibit significant correlation within the same occupation (i.e., product design and market analysis in product manager), as well as for the same skills across different occupations (i.e., product design in backend development engineer and product manager). This demonstrates the necessity of considering all variables as inputs for job skill demand forecasting models, as it captures the interrelationships among variables, preventing the loss of critical information when variables are considered in isolation.
Dataset Limitation. Recent studies [35] suggest that incorporating relationships between different variants can enhance the performance of multivariate time-series forecasting. However, due to the lack of prior knowledge, our Job-SDF dataset does not yet include a graph of relationships between skills, such as predecessor-successor relationships. Instead, we constructed a skill graph based on the co-occurrence of skills in job advertisements from the training data. This graph is included in our dataset. For detailed construction methods, please refer to the Appendix.
4 Benchmark
4.1 Benchmark Models
We evaluated several SOTA time-series learning models using our proposed Job-SDF dataset. These models are categorized into six groups based on their underlying architectures: statistical time series models, RNN-based models, Transformer-based models, MLP-based models, Graph-based models, and Fourier-based models. The implementation details for each model are provided in the Appendix, and the open-source model implementations are available on our GitHub repository.
Statistical Time Series Model. We first consider two statistical methodologies, namely ARIMA [36] and Prophet [37], both of which have been widely used in various contexts. The ARIMA model, which integrates differencing and moving averages within autoregression, has proven effective in forecasting occupational task demands [8]. Prophet decomposes time series data into trend, seasonality, and holiday components, allowing it to handle both linear and nonlinear trends with changepoints. However, these models often struggle to capture complex nonlinear relationships and exhibit suboptimal performance in large-scale data scenarios.
RNN-based Model. RNN-based methods are effective in capturing temporal state transitions through their recurrent structures, making them widely used in various time series forecasting tasks. Notably, LSTM have demonstrated their effectiveness in predicting changes in skill shares over time [9]. However, conventional RNNs often encounter performance degradation when handling excessively long look-back windows and forecast horizons. To address this challenge, SegRNN [12] introduces segment-wise iterations, which reduce the recurrence count within RNNs, thereby significantly enhancing performance in time series forecasting tasks.
Transformer-based Model. Recently, Transformer-based models [38] have gained widespread recognition in long-term time series forecasting due to their global modeling capabilities. Leveraging the attention mechanism, Reformer [17] introduces locally sensitive hashing to approximate attention by grouping similar queries. Informer [13] incorporates low-rank matrices in self-attention mechanisms to accelerate computation. Autoformer[14] employs block decomposition and autocorrelation mechanisms to more effectively capture the intrinsic features of time series data. FedFormer [16] utilizes DFT-based frequency-enhanced attention, obtaining attentive weights through the spectrums of queries and keys and calculating the weighted sum in the frequency domain. To address the challenges of non-stationary time series, the Non-stationary Transformer (NStransformer) [15] introduces a sequence stabilization module and proposes a de-stationary attention mechanism. Additionally, PatchTST [39] is a channel-independent patch time series transformer model that features patching and channel-independence as its key design elements.
MLP-based Model. Multiple Layer Projection (MLP) has been introduced in time series forecasting, demonstrating superior performance compared to transformer-based models in both accuracy and efficiency [18]. Specifically, DLinear [18] uses series decomposition as a pre-processing step before linear regression. FreTS [40] explores a novel approach by applying MLPs in the frequency domain for time series forecasting. TSMixer [19] employs MLPMixer blocks, segments input time series into fixed windows, and applies gated MLP transformations and permutations to enhance accuracy.
Graph-based Models. Graph Neural Networks (GNNs) can learn non-Euclidean relationships, making them effective for identifying associations in structured data and generating joint representations from different perspectives. CHGH [10] uses an adaptive graph enhanced by skill co-occurrence relationships to link skill supply and demand sequences. This fusion of representations across views improves the performance of joint skill supply and demand prediction tasks. Pre-DyGAE [30] targets skill demand prediction from an occupational perspective. It builds an occupation-skill bipartite graph based on the skill demands of occupations and captures the dynamic changes in these relationships. This method allows for predicting both potential occupational skills and skill demands, leveraging a dynamic graph perspective.
Fourier-based Models. By utilizing Fourier projection, FiLM [20] not only captures long-term time dependencies but also effectively reduces noise in forecasting. To address the challenge of non-stationary time-series forecasting, Koopa [21] disentangles time-variant and time-invariant components from complex non-stationary series using a Fourier Filter and designs the Koopman Predictor to forecast dynamics.
4.2 Evaluation Metrics
To evaluate the performance of various benchmark models in job skill demand forecasting tasks, we selected two commonly used regression metrics: MAE and RMSE. MAE is calculated over $H$ observations using the formula: $\frac{1}{H}\sum_{i=1}^{H}|y_{i}-\hat{y}_{i}|$, where $y_{i}$ represents the ground truth value and $\hat{y}_{i}$ is the predicted value. RMSE is calculated as: $\sqrt{\frac{1}{H}\sum_{i=1}^{H}\left(y_{i}-\hat{y}_{i}\right)^{2}}$. Both MAE and RMSE are scale-dependent metrics, which makes them unsuitable for comparison across different granularities. Additionally, these metrics are less sensitive to prediction errors at lower skill demand values. Therefore, we additionally applied SMAPE [22] and RRMSE [41] to assess the performance of various forecasting models. SMAPE considers both the magnitude and direction of errors, making it suitable for comparing forecasts across different scales. RRMSE measures the square root of the average of the squared percentage errors.
$SMAPE=\frac{2}{H}*\sum_{i=1}^{H}\frac{\left|y_{i}-\hat{y}_{i}\right|}{\left|y_{i}\right|+\left|\hat{y}_{i}\right|},\ \ \ RRMSE=\sqrt{\frac{\frac{1}{H}\sum_{i=1}^{H}\left(y_{i}-\hat{y}_{i}\right)^{2}}{\sum_{i=1}^{H}\left(\hat{y}_{i}\right)^{2}}}.$ (3)
Table 1: Performance comparison on MAE and RMSE.
Model L1-Occupation L2-Occupation Region&L1-O Region&L2-O Company
MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE
ARIMA 20.27 256.89 6.46 115.79 3.98 58.65 1.31 27.42 1.31 38.88
Prophet 29.15 356.67 8.95 161.01 5.08 72.21 1.62 33.02 1.55 41.19
LSTM 19.05 194.67 7.09 116.36 3.92 51.59 1.29 23.31 1.35 26.47
SegRNN 12.28 108.28 5.01 68.83 3.14 34.26 1.05 15.96 1.01 16.03
CHGH 22.09 261.49 7.09 116.58 3.91 51.46 1.28 23.24 1.34 26.52
Pre-DyGAE 22.98 187.90 7.04 82.97 4.24 38.62 1.37 17.39 1.24 18.24
Transformer 22.06 215.09 7.58 118.21 4.01 52.04 1.35 23.44 1.26 24.99
Autoformer 23.06 186.76 8.22 100.02 6.45 57.77 2.41 24.10 3.31 38.55
Informer 22.21 205.24 7.43 117.38 3.88 50.13 1.30 23.07 1.26 24.92
Reformer 22.11 204.35 7.46 116.60 3.91 50.95 1.25 22.81 1.54 27.37
FEDformer 22.87 181.93 7.46 88.97 4.63 43.21 1.98 21.73 2.43 26.92
NStransformer 17.36 149.46 5.75 86.24 3.45 37.09 1.15 17.45 2.13 34.83
PatchTST 14.91 141.06 5.15 78.86 3.10 35.38 1.04 16.57 1.01 19.09
DLinear 16.61 154.88 5.44 81.61 3.24 36.67 1.07 16.79 1.05 18.85
TSMixer 21.34 192.85 8.14 106.65 5.81 62.14 5.95 68.26 13.96 144.96
FreTS 16.47 167.61 6.52 106.39 3.65 47.81 1.22 21.92 1.26 25.39
FiLM 12.95 117.17 5.08 65.65 3.24 29.90 1.14 14.01 1.17 15.87
Koopa 19.91 179.30 6.05 91.87 3.53 40.73 1.15 18.71 1.08 20.18
4.3 Benchmark Results
Overall Performance. In Table 1, we present the performance of various models evaluated using two metrics: MAE and RMSE. The following conclusions can be drawn: (1) The traditional statistical method, Prophet, demonstrates relatively poor predictive performance. This may be due to seasonal and holiday factors not being the primary influencers in skill demand prediction. (2) Most Transformer-based models, including Transformer, Autoformer, Informer, and Reformer, exhibit subpar overall predictive performance. This is likely because these models are designed to address long-range temporal dependencies, which are not well-suited for the current shorter time series context. (3) In contrast, PatchTST, unlike these Transformer-based models that perform point-wise modeling of time series, segments the time series into patches and inputs them into the Transformer. This allows the model to focus on more local information. A similar idea is also employed in the SegRNN. This strategy significantly enhances the performance of these models in predicting job skill demand. (4) The performance of different linear models on our dataset varies significantly. For instance, DLinear outperforms most Transformer-based models, while TSMixer performs poorly. This discrepancy may be due to the tendency of more complex MLP-based models to overfit our dataset. (5) CHGH and Pre-DyGAE exhibit poor performance in the separate skill demand forecasting scenario, likely due to a mismatch between their model design and the context of our dataset. Specifically, CHGH relies on sequential data from the supply side of skills, which is lacking in our dataset. Conversely, Pre-DyGAE focuses more on predicting whether a skill will be required by an occupation in the future. (6) Finally, FiLM achieved the best performance in most cases, demonstrating the robustness of the denoising-based model.
Low-Demand Skill Prediction Performance. Considering the varying nature of skill demand values, we further employed SMAPE and RRMSE metrics to focus on the predictive performance of different models for low-demand skills. As shown in Table 2, the experimental results indicate the following: (1) PatchTST achieved the best SMAPE performance in most cases, validating its ability to more accurately predict the trends of low-demand skills. (2) Based on scale-independent metrics, we can compare the performance of models at different granularities. It can be observed that RRMSE exhibits a significant trend of variation across different granularities; specifically, as the granularity becomes finer, the RRMSE performance deteriorates. This indicates that predicting skill demand at finer granularities is more challenging. Additionally, FiLM shows the least variation across multiple granularities, further validating its ability to provide stable and reliable predictions under varying granularities and demand value ranges. (3) Although Koopa performs averagely on MAE and RMSE metrics, it excels in predicting low-demand skills, particularly in terms of SMAPE. Similarly, NStransformer also performs well in scenarios focusing on low-demand skill predictions. This success can be attributed to both methods being designed to handle non-stationary time series. They effectively filter noise from historical sequences and restore intrinsic non-stationary information into time-dependent relationships, making them more adept at handling the fluctuating nature of low-demand skill time series data.
Table 2: Performance comparison on SMAPE and RRMSE.
Model L1-Occupation (%) L2-Occupation (%) Region&L1-O (%) Region&L2-O (%) Company (%)
SMAPE RRMSE SMAPE RRMSE SMAPE RRMSE SMAPE RRMSE SMAPE RRMSE
ARIMA 35.72 47.89 25.00 58.87 23.86 58.07 13.58 73.57 20.17 147.94
Prophet 41.22 67.78 28.35 88.47 26.75 71.60 15.07 93.04 22.31 167.77
LSTM 41.38 57.90 32.85 83.70 31.58 68.40 22.93 87.36 30.26 174.40
SegRNN 39.81 37.58 33.35 50.53 35.30 48.53 23.84 61.90 33.07 86.27
CHGH 40.27 66.05 29.60 84.10 28.11 68.42 17.42 87.45 26.72 176.70
PreDyGAE 49.87 83.67 60.54 83.60 59.32 66.56 72.67 98.09 26.21 145.73
Transformer 55.59 64.25 44.23 84.27 31.15 76.16 33.04 86.87 27.61 164.36
Autoformer 70.28 53.75 74.37 63.40 90.14 65.57 91.51 74.46 107.05 99.60
Informer 56.85 58.18 44.04 88.72 34.75 69.59 29.29 90.15 32.41 164.37
Reformer 56.58 61.35 40.58 83.70 32.21 72.87 20.86 90.85 45.25 169.87
FEDformer 69.30 54.03 69.29 60.00 73.17 52.69 81.73 70.06 94.19 97.97
NStransformer 38.11 47.19 26.30 60.73 24.98 48.89 14.55 63.29 24.20 100.78
PatchTST 34.70 51.17 24.52 58.80 25.15 44.96 13.50 67.48 19.89 115.34
DLinear 41.84 52.89 34.35 60.22 33.47 51.05 25.77 64.65 30.71 108.66
TSMixer 56.59 61.17 72.29 99.35 82.48 87.29 120.85 96.49 155.20 102.14
FreTS 39.76 54.42 30.18 80.44 28.58 66.11 17.62 85.04 27.24 174.56
FiLM 39.51 37.55 29.65 43.86 28.79 37.66 17.24 47.75 25.72 76.92
Koopa 37.84 58.30 25.72 65.34 24.41 57.81 13.98 74.00 20.43 123.96
Table 3: Performance comparison on data with structural breaks on MAE and RMSE.
Model L1-Occupation L2-Occupation Region&L1-O Region&L2-O Company
MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE
LSTM 87.30 554.46 57.95 400.22 18.99 149.53 7.91 52.38 24.40 159.02
SegRNN 61.92 390.54 43.97 276.57 15.85 114.04 6.56 37.84 17.98 112.13
CHGH 94.30 629.32 58.06 401.45 19.00 149.75 7.90 52.50 24.37 159.44
PreGyGAE 78.35 493.83 48.69 336.15 17.49 136.66 7.31 38.88 19.76 164.43
Transformer 98.66 580.58 61.73 404.17 19.37 151.12 8.45 55.46 22.41 152.27
Autoformer 107.22 533.06 67.66 350.97 26.84 156.50 12.19 63.04 44.10 208.96
Informer 98.89 570.35 59.95 402.75 19.03 146.91 7.72 49.15 22.37 151.87
Reformer 98.14 569.83 60.71 401.21 19.25 149.91 7.52 49.10 25.65 160.69
FEDformer 105.43 532.24 62.10 325.10 20.49 128.45 10.37 55.47 34.09 155.28
NStransformer 82.43 462.24 49.30 318.44 16.59 119.91 6.85 37.56 40.05 196.03
PatchTST 77.44 474.86 45.02 303.76 14.88 111.01 6.56 38.60 18.03 127.72
DLinear 81.17 485.25 46.67 307.34 15.94 118.94 6.50 37.72 18.18 124.32
TSMixer 107.47 614.93 83.60 479.39 29.99 187.08 25.83 190.29 155.10 766.58
FreTS 82.45 537.12 56.54 393.38 18.55 148.33 7.88 52.87 24.21 160.01
FiLM 62.86 404.82 42.63 260.99 14.31 101.23 6.37 32.28 18.78 110.65
Koopa 91.26 516.75 50.44 324.15 17.43 128.39 7.07 41.29 19.04 133.26
Table 4: Performance comparison on data with structural breaks on RRMSE and SMAPE.
Model L1-Occupation (%) L2-Occupation (%) Region&L1-O (%) Region&L2-O (%) Company (%)
SMAPE RRMSE SMAPE RRMSE SMAPE RRMSE SMAPE RRMSE SMAPE RRMSE
LSTM 43.78 58.05 48.93 84.46 46.64 78.31 42.03 58.48 68.38 187.30
SegRNN 39.22 37.80 43.09 51.14 45.17 54.31 39.41 41.40 57.45 89.65
CHGH 44.91 66.31 48.90 84.87 45.43 78.32 39.79 58.89 68.36 189.91
PreDyGAE 52.35 47.15 56.56 59.31 52.06 61.22 44.13 42.31 70.26 106.88
Transformer 50.01 64.47 53.10 84.95 46.50 86.56 47.67 61.23 64.92 177.43
Autoformer 63.46 54.08 68.62 64.14 87.93 68.97 88.95 63.85 115.00 100.60
Informer 51.11 58.40 51.89 89.70 47.81 80.86 44.90 57.55 65.11 177.16
Reformer 50.79 61.59 51.51 84.53 46.86 84.15 40.81 58.59 72.36 181.36
FEDformer 62.83 54.37 64.37 60.84 72.24 58.55 80.03 54.29 103.27 100.65
NStransformer 45.36 47.46 47.63 61.85 43.04 57.60 36.72 39.72 170.57 113.87
PatchTST 40.89 51.48 43.26 59.69 41.51 51.85 34.74 43.12 55.26 122.56
DLinear 43.14 53.20 45.25 61.13 45.26 58.80 41.15 41.71 57.65 115.24
TSMixer 54.31 61.31 76.08 99.84 85.12 95.81 117.39 93.66 160.55 102.23
FreTS 42.44 54.59 48.24 81.17 45.39 75.43 39.85 57.83 68.39 187.94
FiLM 38.96 37.82 44.23 44.52 44.95 43.06 40.05 30.80 56.37 80.77
Koopa 46.45 58.59 47.13 66.28 42.60 66.20 36.24 47.48 58.98 131.77
Performance on Skill Demand Series with Structural Breaks. As described in Section 3.3, in the dynamically changing job market, skill demand time series data exhibit structural breaks. To assess the impact of this phenomenon on different models in the skill demand forecasting task, we used the Chow test to detect structural breaks in the skill demand time series. The corresponding predictive performance of different models is presented in Tables 3 and 4. We observe the following phenomena: (1) Compared to the predictive performance on the full dataset, the performance on time series data with structural breaks is significantly worse. This finding underscores the complexity and unpredictability of skill trends that experience structural breaks. (2) FiLM has achieved results close to the overall skill demand prediction in terms of SMAPE and RRMSE metrics. This validates that FiLM can effectively mitigate the disruptive impact of structural breaks on skill demand forecasting. (3) Furthermore, while the overall predictive performance of skill demand forecasting at both the Region&L2-O and Company granularity levels is similar, significant differences emerge when forecasting skills experiencing structural breaks. This suggests that skills undergoing structural breaks display more predictable patterns at the Region&L2-O granularity level compared to the Company level, making them relatively easier to forecast.
5 Conclusion
In this work, we introduced Job-SDF, a dataset designed for training and benchmarking job-skill demand forecasting models. Compiled from 10.35 million public job advertisements collected from major online recruitment platforms in China between January 2021 and December 2023, this dataset includes monthly recruitment demand for 2,324 types of skills across 52 occupations, 521 companies, and 7 regions. Using this dataset, we validated a wide range of time-series forecasting approaches, including statistical models, RNN-based models, Transformer-based models, MLP-based models, Graph-based models, and Fourier-based models. Furthermore, we conducted extensive experiments to compare the performance of various methods in predicting skill demand at different granularities. We hope that Job-SDF will facilitate further research in this field.
Appendix A Computational Resource
Due to inherent design and size constraints of the models combined with varying data sizes at different granularities, the deployment environments for each model are distinct. The CHGH model, which requires over 80GB of memory, is exclusively deployed on CPU platforms to accommodate its substantial resource demands. In contrast, the PreDyGAE model operates solely on GPU infrastructure, leveraging the computational efficiencies of the NVIDIA A800 GPUs. For other models, deployment strategies are tailored according to the granularity of the data. Experiments at the labor market, regions, L1 occupations, L2 occupations, and Region & L1 occupations granularities are conducted on GPUs, capitalizing on the enhanced processing capabilities of these units for handling moderate data volumes. However, at the granularities of Region & L2 and company, where data volumes are significantly larger, deployment shifts to CPUs. Overall, the training time of different models are shown in Table S5.
Table S5: Training time (minute) of different models for job skill demand forecsting.
Model Market Region L1-O L2-O R&L1-O R&L2-O Company
LSTM 0-0.5 0-0.5 0-0.5 0-0.5 0-0.5 37.7 39.0
SegRNN 0-0.5 0-0.5 0-0.5 0-0.5 0-0.5 342.8 458.2
CHGH 17.7 132.8 170.2 258.3 1300.3 490.6 6604.2
PreDyGAE 1-10 16.5 30.0 48.1 48.1 88.2 126.2
Transformer 0-0.5 0-0.5 0-0.5 0.5-1 0.5-1 128.2 166.5
Autoformer 0-0.5 0-0.5 0-0.5 0-0.5 0.5-1 304.3 325.0
Informer 0-0.5 0-0.5 0-0.5 0-0.5 0-0.5 133.8 171.7
Reformer 0-0.5 0-0.5 0-0.5 0-0.5 0-0.5 36.0 52.5
FEDformer 0-0.5 0-0.5 0-0.5 0-0.5 0-0.5 193.8 198.5
NStransformer 0-0.5 0.5-1 0.5-1 0-0.5 0.5-1 128.5 195.7
PatchTST 0-0.5 0-0.5 0-0.5 0-0.5 0.5-1 1202.8 2558.0
DLinear 0-0.5 0-0.5 0.5-1 0-0.5 0-0.5 20.0 39.1
TSMixer 0-0.5 0-0.5 0-0.5 0-0.5 0-0.5 24.0 97.0
FreTS 0-0.5 0-0.5 0-0.5 0-0.5 0-0.5 85.0 200.0
FiLM 0-0.5 0.5-1 0.5-1 1-10 1-10 598.0 1464.7
Koopa 0-0.5 0-0.5 0-0.5 0-0.5 0-0.5 38.3 68.5
Appendix B Additional Experimental Results
Due to page limitations in the main text, we have included additional experimental content in the appendix. First, we present the results of repeated trials of the benchmark models discussed in the main text in the first subsection. Subsequently, we focus on the performance of existing benchmark models in predicting demand for low-frequency skills. Further, we have constructed a co-occurrence relationship between skills as prior knowledge based on the training set and employed various Graph Neural Network (GNN)-based multivariate time series forecasting methods in the task of job skill demand forecasting, demonstrating promising results. Finally, considering that the skill demand proportion may be more meaningful than the skill demand volume in certain contexts, we have constructed a dataset for skill demand proportion and showcased the performance of benchmark models on this task.
B.1 Repeated Experiments on Job Skill Demand Forecasting
To demonstrate the robustness and reliability of our experimental results, we first repeated the experiments multiple times as described in the main text. Additionally, we extended our analysis to include experiments across the entire labor market and at various regional granularities.
Implementment Details.
We utilized the Time-series-Library  [^2] https://github.com/thuml/Time-Series-Library to implement some of the models. The hyperparameters were uniformly set as follows: a learning rate of 0.0001, 20 epochs of training, a hidden layer dimension of 2048, the GELU activation function, and MSE loss as the loss function. Early stopping was employed to prevent overfitting by terminating training early when necessary. Data from the first 24 months were used for pre-training, and the model was fine-tuned using the next 6 months to capture trend changes. Finally, the model was used to infer skill demands for the last 6 months. All other hyperparameters were kept consistent with those in the original paper. To ensure the reliability of our findings, we repeated these experiments four times, using random seeds set to 0, 1, 2, and 3, respectively.
Overall Performance
Table S6 displays the mean and standard deviation results of repeated experiments on the benchmark models for the job skill demand forecasting task as presented in the main text. Initially, we supplement the experimental results at the overall labor market level and regional granularity, where the RMSE averages over 1000. In cases of coarser granularity, due to the larger base of demand values, the prediction deviations are significant.
Table S6: Overall performance comparisons on repeated experiments.
Model Market Region L1-O L2-O R&L1-O R&L2-O Company
MAE LSTM $314.54\pm_{0.57}$ $49.92\pm_{0.0}$ $24.43\pm_{4.91}$ $8.04\pm_{0.87}$ $4.44\pm_{0.47}$ $1.45\pm_{0.15}$ $1.49\pm_{0.13}$
SegRNN $190.05\pm_{0.37}$ $35.92\pm_{0.47}$ $16.37\pm_{3.73}$ $5.81\pm_{0.73}$ $3.68\pm_{0.5}$ $1.32\pm_{0.25}$ $1.23\pm_{0.2}$
CHGH $315.47\pm_{0.04}$ $50.03\pm_{0.02}$ $25.62\pm_{3.23}$ $8.04\pm_{0.87}$ $4.43\pm_{0.48}$ $1.51\pm_{0.7}$ $1.54\pm_{0.73}$
PreDyGAE $189.95\pm_{0.01}$ $35.84\pm_{0.09}$ $21.72\pm_{1.15}$ $6.81\pm_{0.21}$ $4.08\pm_{0.15}$ $1.85\pm_{0.43}$ $1.85\pm_{0.56}$
Transformer $340.72\pm_{0.79}$ $54.98\pm_{0.13}$ $27.43\pm_{4.9}$ $8.93\pm_{1.24}$ $4.92\pm_{0.83}$ $1.75\pm_{0.36}$ $1.69\pm_{0.39}$
Autoformer $465.86\pm_{2.78}$ $60.9\pm_{0.64}$ $31.97\pm_{8.13}$ $9.97\pm_{1.6}$ $6.68\pm_{0.21}$ $2.6\pm_{0.17}$ $2.85\pm_{0.42}$
Informer $340.76\pm_{1.6}$ $55.08\pm_{0.35}$ $27.54\pm_{4.87}$ $8.81\pm_{1.26}$ $4.87\pm_{0.91}$ $1.73\pm_{0.39}$ $1.69\pm_{0.39}$
Reformer $344.9\pm_{2.62}$ $54.83\pm_{0.2}$ $27.35\pm_{4.79}$ $8.88\pm_{1.3}$ $4.9\pm_{0.9}$ $1.71\pm_{0.42}$ $1.8\pm_{0.24}$
FEDformer $469.46\pm_{4.52}$ $60.37\pm_{0.19}$ $31.87\pm_{8.22}$ $9.54\pm_{1.9}$ $5.83\pm_{1.1}$ $2.41\pm_{0.39}$ $2.52\pm_{0.08}$
NStransformer $208.23\pm_{3.43}$ $37.39\pm_{0.46}$ $19.13\pm_{1.61}$ $6.29\pm_{0.5}$ $3.79\pm_{0.31}$ $1.26\pm_{0.1}$ $1.64\pm_{0.44}$
PatchTST $204.94\pm_{12.41}$ $36.33\pm_{2.23}$ $18.69\pm_{3.45}$ $6.12\pm_{0.89}$ $3.69\pm_{0.54}$ $1.23\pm_{0.18}$ $1.21\pm_{0.18}$
DLinear $201.05\pm_{1.95}$ $35.45\pm_{0.56}$ $18.23\pm_{1.48}$ $5.89\pm_{0.41}$ $3.55\pm_{0.28}$ $1.21\pm_{0.12}$ $1.18\pm_{0.11}$
TSMixer $517.95\pm_{1.49}$ $67.56\pm_{11.06}$ $30.83\pm_{8.66}$ $10.54\pm_{2.19}$ $5.95\pm_{0.12}$ $3.58\pm_{2.16}$ $6.75\pm_{6.58}$
FreTS $310.62\pm_{4.57}$ $49.52\pm_{0.73}$ $23.1\pm_{6.05}$ $7.74\pm_{1.11}$ $4.3\pm_{0.59}$ $1.41\pm_{0.17}$ $1.45\pm_{0.17}$
FiLM $201.7\pm_{24.86}$ $36.18\pm_{2.29}$ $16.79\pm_{3.51}$ $5.9\pm_{0.75}$ $3.7\pm_{0.42}$ $1.28\pm_{0.13}$ $1.31\pm_{0.12}$
Koopa $205.12\pm_{4.95}$ $35.95\pm_{0.5}$ $19.9\pm_{0.01}$ $6.19\pm_{0.13}$ $3.68\pm_{0.13}$ $1.2\pm_{0.05}$ $1.15\pm_{0.06}$
RMSE LSTM $1799.7\pm_{0.49}$ $341.97\pm_{0.07}$ $308.58\pm_{103.99}$ $148.32\pm_{29.18}$ $66.87\pm_{13.95}$ $30.01\pm_{6.11}$ $35.35\pm_{8.1}$
SegRNN $941.59\pm_{1.08}$ $194.73\pm_{1.59}$ $154.02\pm_{41.75}$ $78.84\pm_{9.14}$ $38.6\pm_{3.97}$ $18.21\pm_{2.05}$ $21.05\pm_{4.58}$
CHGH $1803.56\pm_{0.73}$ $344.15\pm_{0.11}$ $335.97\pm_{67.99}$ $148.94\pm_{29.54}$ $67.12\pm_{14.29}$ $19.3\pm_{12.73}$ $20.61\pm_{14.53}$
PreDyGAE $925.27\pm_{0.01}$ $193.94\pm_{0.04}$ $186.36\pm_{1.4}$ $84.55\pm_{1.44}$ $40.37\pm_{1.6}$ $18.79\pm_{1.27}$ $22.05\pm_{3.48}$
Transformer $1829.87\pm_{2.96}$ $360.87\pm_{6.13}$ $318.17\pm_{94.1}$ $151.16\pm_{30.08}$ $67.66\pm_{14.26}$ $30.42\pm_{6.37}$ $35.12\pm_{9.25}$
Autoformer $2177.05\pm_{87.58}$ $313.84\pm_{8.55}$ $270.26\pm_{76.23}$ $122.29\pm_{20.33}$ $62.56\pm_{4.38}$ $26.57\pm_{2.26}$ $35.47\pm_{2.81}$
Informer $1854.81\pm_{3.0}$ $359.97\pm_{7.72}$ $314.97\pm_{100.17}$ $148.73\pm_{28.62}$ $67.01\pm_{15.41}$ $30.14\pm_{6.45}$ $35.1\pm_{9.3}$
Reformer $1890.32\pm_{13.72}$ $357.61\pm_{1.23}$ $314.03\pm_{100.12}$ $152.43\pm_{32.71}$ $67.58\pm_{15.18}$ $30.21\pm_{6.76}$ $36.14\pm_{8.01}$
FEDformer $2137.44\pm_{58.69}$ $321.53\pm_{2.51}$ $268.73\pm_{79.24}$ $113.47\pm_{22.37}$ $56.36\pm_{12.0}$ $25.64\pm_{3.57}$ $30.9\pm_{3.64}$
NStransformer $1121.07\pm_{25.29}$ $236.56\pm_{12.38}$ $198.03\pm_{44.33}$ $103.2\pm_{15.48}$ $45.99\pm_{8.13}$ $21.42\pm_{3.62}$ $32.8\pm_{1.86}$
PatchTST $1098.82\pm_{43.54}$ $220.21\pm_{7.5}$ $204.56\pm_{57.97}$ $97.9\pm_{17.38}$ $45.32\pm_{9.07}$ $20.88\pm_{3.94}$ $25.48\pm_{5.83}$
DLinear $1107.0\pm_{28.41}$ $222.14\pm_{4.31}$ $211.02\pm_{51.25}$ $99.81\pm_{16.61}$ $46.04\pm_{8.55}$ $21.12\pm_{3.95}$ $25.7\pm_{6.25}$
TSMixer $2578.82\pm_{27.27}$ $438.06\pm_{67.68}$ $350.22\pm_{143.66}$ $172.63\pm_{60.23}$ $78.88\pm_{15.28}$ $48.76\pm_{17.8}$ $82.91\pm_{56.64}$
FreTS $1763.69\pm_{25.03}$ $336.65\pm_{4.04}$ $289.29\pm_{111.07}$ $140.54\pm_{31.17}$ $63.72\pm_{14.52}$ $28.68\pm_{6.17}$ $34.0\pm_{7.86}$
FiLM $1071.14\pm_{223.85}$ $216.77\pm_{38.28}$ $173.2\pm_{51.15}$ $83.31\pm_{16.12}$ $39.26\pm_{8.54}$ $18.23\pm_{3.85}$ $22.53\pm_{6.08}$
Koopa $1147.88\pm_{31.89}$ $228.36\pm_{7.6}$ $227.46\pm_{43.96}$ $106.03\pm_{12.93}$ $49.37\pm_{7.89}$ $22.53\pm_{3.49}$ $27.06\pm_{6.28}$
SMAPE(%) LSTM $42.33\pm_{0.04}$ $49.78\pm_{0.03}$ $42.31\pm_{0.85}$ $33.14\pm_{0.27}$ $31.79\pm_{0.19}$ $22.67\pm_{0.24}$ $30.42\pm_{0.14}$
SegRNN $39.67\pm_{0.26}$ $51.14\pm_{0.53}$ $42.16\pm_{2.14}$ $34.62\pm_{1.16}$ $36.02\pm_{0.65}$ $30.62\pm_{6.19}$ $32.77\pm_{0.27}$
CHGH $42.21\pm_{0.01}$ $49.03\pm_{0.03}$ $40.83\pm_{0.51}$ $29.98\pm_{0.35}$ $28.42\pm_{0.28}$ $46.97\pm_{9.54}$ $30.69\pm_{14.64}$
PreDyGAE $42.1\pm_{0.14}$ $64.37\pm_{1.24}$ $75.3\pm_{23.21}$ $59.69\pm_{0.77}$ $46.51\pm_{11.69}$ $76.6\pm_{3.58}$ $60.36\pm_{31.18}$
Transformer $49.86\pm_{0.46}$ $60.59\pm_{0.02}$ $58.06\pm_{2.26}$ $52.7\pm_{7.73}$ $47.37\pm_{14.81}$ $47.03\pm_{12.77}$ $45.69\pm_{16.5}$
Autoformer $73.47\pm_{0.36}$ $79.08\pm_{0.19}$ $79.44\pm_{8.36}$ $79.16\pm_{4.38}$ $86.16\pm_{3.64}$ $86.32\pm_{4.73}$ $91.12\pm_{14.54}$
Informer $49.2\pm_{0.2}$ $60.45\pm_{0.03}$ $58.59\pm_{1.59}$ $52.61\pm_{7.82}$ $48.84\pm_{12.86}$ $45.52\pm_{14.82}$ $47.61\pm_{13.88}$
Reformer $48.69\pm_{0.17}$ $60.48\pm_{0.17}$ $58.49\pm_{1.74}$ $51.26\pm_{9.75}$ $47.8\pm_{14.24}$ $42.15\pm_{19.43}$ $52.76\pm_{6.85}$
FEDformer $73.57\pm_{0.53}$ $78.11\pm_{0.55}$ $79.26\pm_{9.09}$ $76.98\pm_{7.02}$ $78.54\pm_{4.9}$ $82.05\pm_{0.3}$ $86.29\pm_{7.21}$
NStransformer $38.88\pm_{0.09}$ $47.74\pm_{0.14}$ $38.94\pm_{0.76}$ $27.49\pm_{1.08}$ $26.43\pm_{1.33}$ $15.47\pm_{0.84}$ $23.75\pm_{0.41}$
PatchTST $37.56\pm_{2.15}$ $46.52\pm_{2.49}$ $37.82\pm_{2.85}$ $26.98\pm_{2.25}$ $26.7\pm_{1.42}$ $15.2\pm_{1.55}$ $22.29\pm_{2.19}$
DLinear $37.7\pm_{1.82}$ $49.27\pm_{3.85}$ $41.79\pm_{0.05}$ $34.65\pm_{0.28}$ $34.07\pm_{0.54}$ $26.6\pm_{0.75}$ $31.42\pm_{0.65}$
TSMixer $73.63\pm_{1.83}$ $72.27\pm_{9.41}$ $61.01\pm_{4.04}$ $68.37\pm_{3.58}$ $65.19\pm_{15.78}$ $79.28\pm_{37.95}$ $95.62\pm_{54.39}$
FreTS $42.81\pm_{0.56}$ $50.46\pm_{0.62}$ $41.53\pm_{1.62}$ $31.26\pm_{0.98}$ $29.76\pm_{1.08}$ $18.89\pm_{1.16}$ $28.37\pm_{1.03}$
FiLM $37.79\pm_{1.75}$ $46.85\pm_{4.01}$ $41.03\pm_{1.38}$ $30.26\pm_{0.56}$ $29.49\pm_{0.64}$ $17.65\pm_{0.37}$ $26.19\pm_{0.42}$
Koopa $36.26\pm_{0.39}$ $44.7\pm_{0.95}$ $36.69\pm_{1.05}$ $25.59\pm_{0.12}$ $24.51\pm_{0.09}$ $14.13\pm_{0.14}$ $20.62\pm_{0.18}$
RRMSE(%) LSTM $52.7\pm_{0.09}$ $52.87\pm_{0.03}$ $73.45\pm_{14.2}$ $97.42\pm_{12.52}$ $80.24\pm_{10.81}$ $102.19\pm_{13.54}$ $210.9\pm_{33.32}$
SegRNN $20.55\pm_{0.05}$ $22.49\pm_{0.39}$ $32.2\pm_{4.91}$ $41.44\pm_{8.3}$ $39.72\pm_{8.04}$ $49.27\pm_{11.53}$ $71.42\pm_{13.56}$
CHGH $52.92\pm_{0.03}$ $53.46\pm_{0.02}$ $77.26\pm_{10.23}$ $98.13\pm_{12.8}$ $80.86\pm_{11.35}$ $54.98\pm_{47.9}$ $100.68\pm_{96.78}$
PreDyGAE $20.07\pm_{0.0}$ $22.65\pm_{0.01}$ $51.16\pm_{29.68}$ $54.87\pm_{26.22}$ $47.52\pm_{17.38}$ $64.98\pm_{30.23}$ $96.4\pm_{45.03}$
Transformer $52.49\pm_{0.18}$ $55.57\pm_{1.14}$ $75.98\pm_{10.71}$ $98.09\pm_{12.61}$ $82.63\pm_{5.9}$ $101.34\pm_{13.21}$ $200.07\pm_{32.6}$
Autoformer $49.81\pm_{1.73}$ $38.47\pm_{1.72}$ $54.53\pm_{0.71}$ $64.59\pm_{1.09}$ $61.08\pm_{4.1}$ $71.01\pm_{3.15}$ $103.89\pm_{3.92}$
Informer $54.11\pm_{0.2}$ $55.59\pm_{2.17}$ $71.42\pm_{12.08}$ $97.76\pm_{8.25}$ $81.04\pm_{10.45}$ $102.15\pm_{10.95}$ $198.51\pm_{31.17}$
Reformer $55.75\pm_{0.54}$ $54.83\pm_{0.62}$ $76.12\pm_{13.48}$ $98.47\pm_{13.48}$ $81.65\pm_{8.01}$ $103.29\pm_{11.35}$ $206.04\pm_{33.02}$
FEDformer $48.25\pm_{2.89}$ $39.48\pm_{0.85}$ $56.52\pm_{2.27}$ $60.81\pm_{0.74}$ $55.93\pm_{2.95}$ $69.07\pm_{0.9}$ $102.11\pm_{3.78}$
NStransformer $26.75\pm_{0.76}$ $30.42\pm_{1.94}$ $43.64\pm_{3.24}$ $59.15\pm_{1.44}$ $49.08\pm_{0.17}$ $62.42\pm_{0.79}$ $107.45\pm_{6.09}$
PatchTST $26.38\pm_{1.12}$ $28.31\pm_{1.01}$ $47.51\pm_{3.34}$ $56.73\pm_{1.89}$ $47.96\pm_{2.74}$ $64.14\pm_{3.05}$ $108.92\pm_{5.86}$
DLinear $26.6\pm_{1.35}$ $28.58\pm_{1.32}$ $49.04\pm_{3.52}$ $58.69\pm_{1.4}$ $51.47\pm_{0.38}$ $64.72\pm_{0.06}$ $110.8\pm_{1.95}$
TSMixer $74.69\pm_{3.7}$ $69.2\pm_{15.69}$ $86.56\pm_{23.18}$ $118.56\pm_{17.54}$ $99.17\pm_{10.85}$ $100.88\pm_{4.01}$ $136.0\pm_{30.91}$
FreTS $51.01\pm_{0.94}$ $51.53\pm_{0.82}$ $69.56\pm_{13.82}$ $92.39\pm_{10.91}$ $76.55\pm_{9.53}$ $97.37\pm_{11.25}$ $200.81\pm_{23.97}$
FiLM $25.72\pm_{7.38}$ $27.83\pm_{7.19}$ $36.13\pm_{1.29}$ $43.32\pm_{0.5}$ $38.93\pm_{1.16}$ $48.51\pm_{0.7}$ $79.66\pm_{2.5}$
Koopa $27.81\pm_{1.3}$ $29.66\pm_{1.55}$ $52.74\pm_{5.07}$ $62.29\pm_{2.78}$ $56.41\pm_{1.28}$ $70.91\pm_{2.82}$ $122.3\pm_{1.52}$
Performance on Skill Demand Series with Structural Breaks
Table S7: Performance comparisons on skill demand series with structural breaks.
Model Market Region L1-O L2-O R&L1-O R&L2-O Company
MAE LSTM $423.99\pm_{0.56}$ $109.63\pm_{0.04}$ $101.68\pm_{13.13}$ $63.39\pm_{4.97}$ $21.02\pm_{1.85}$ $8.55\pm_{0.59}$ $26.23\pm_{1.67}$
SegRNN $256.67\pm_{0.54}$ $76.09\pm_{1.01}$ $68.08\pm_{5.62}$ $44.93\pm_{0.88}$ $16.12\pm_{0.24}$ $7.07\pm_{0.47}$ $18.91\pm_{0.85}$
CHGH $425.17\pm_{0.04}$ $109.82\pm_{0.06}$ $104.41\pm_{9.23}$ $63.54\pm_{5.0}$ $18.11\pm_{1.93}$ $15.16\pm_{4.33}$ $19.75\pm_{13.35}$
PreDyGAE $296.27\pm_{0.01}$ $85.05\pm_{5.06}$ $84.11\pm_{8.87}$ $56.28\pm_{7.2}$ $16.57\pm_{0.84}$ $9.51\pm_{0.19}$ $22.9\pm_{0.13}$
Transformer $460.45\pm_{2.31}$ $120.62\pm_{0.4}$ $113.11\pm_{13.19}$ $69.45\pm_{7.05}$ $22.4\pm_{2.76}$ $9.42\pm_{0.88}$ $26.93\pm_{4.12}$
Autoformer $632.29\pm_{2.76}$ $131.92\pm_{1.52}$ $131.01\pm_{21.72}$ $75.4\pm_{7.07}$ $25.82\pm_{0.93}$ $11.76\pm_{0.39}$ $36.65\pm_{6.8}$
Informer $462.88\pm_{1.99}$ $120.89\pm_{0.68}$ $113.36\pm_{13.21}$ $68.13\pm_{7.46}$ $22.22\pm_{2.91}$ $9.1\pm_{1.26}$ $26.82\pm_{4.06}$
Reformer $465.72\pm_{3.72}$ $120.58\pm_{0.34}$ $112.51\pm_{13.12}$ $69.02\pm_{7.59}$ $22.4\pm_{2.88}$ $9.09\pm_{1.44}$ $28.16\pm_{2.29}$
FEDformer $645.02\pm_{6.45}$ $131.05\pm_{0.27}$ $130.18\pm_{22.59}$ $72.12\pm_{9.14}$ $23.06\pm_{2.34}$ $11.01\pm_{0.59}$ $32.89\pm_{1.1}$
NStransformer $285.18\pm_{6.1}$ $81.54\pm_{1.2}$ $80.2\pm_{2.03}$ $49.66\pm_{0.33}$ $16.88\pm_{0.26}$ $7.07\pm_{0.2}$ $28.34\pm_{10.69}$
PatchTST $282.34\pm_{17.19}$ $79.1\pm_{4.75}$ $80.46\pm_{2.76}$ $48.45\pm_{3.13}$ $16.29\pm_{1.29}$ $7.01\pm_{0.41}$ $19.65\pm_{1.48}$
DLinear $277.33\pm_{2.32}$ $77.47\pm_{0.99}$ $77.64\pm_{3.23}$ $46.5\pm_{0.16}$ $16.03\pm_{0.08}$ $6.64\pm_{0.12}$ $18.73\pm_{0.51}$
TSMixer $698.27\pm_{6.79}$ $146.22\pm_{23.62}$ $130.91\pm_{21.4}$ $88.39\pm_{4.38}$ $28.49\pm_{1.37}$ $16.46\pm_{8.55}$ $79.02\pm_{69.45}$
FreTS $419.07\pm_{6.85}$ $108.63\pm_{1.75}$ $98.46\pm_{14.62}$ $62.17\pm_{5.14}$ $20.54\pm_{1.82}$ $8.47\pm_{0.54}$ $25.89\pm_{1.53}$
FiLM $277.66\pm_{34.84}$ $78.58\pm_{6.24}$ $69.56\pm_{6.12}$ $45.37\pm_{2.5}$ $15.26\pm_{0.87}$ $6.79\pm_{0.39}$ $19.86\pm_{0.98}$
Koopa $282.85\pm_{6.89}$ $78.75\pm_{1.42}$ $83.25\pm_{7.31}$ $48.75\pm_{1.55}$ $16.92\pm_{0.47}$ $6.93\pm_{0.13}$ $19.16\pm_{0.11}$
RMSE LSTM $2148.33\pm_{0.79}$ $534.61\pm_{0.67}$ $704.41\pm_{136.89}$ $466.21\pm_{60.24}$ $178.29\pm_{26.25}$ $58.54\pm_{5.62}$ $192.72\pm_{30.76}$
SegRNN $1130.19\pm_{0.81}$ $304.2\pm_{2.76}$ $387.1\pm_{3.14}$ $264.53\pm_{10.99}$ $110.27\pm_{3.44}$ $36.38\pm_{1.34}$ $120.84\pm_{7.95}$
CHGH $2153.77\pm_{0.87}$ $537.97\pm_{0.18}$ $735.74\pm_{97.15}$ $468.27\pm_{61.0}$ $179.17\pm_{26.85}$ $41.0\pm_{28.76}$ $203.78\pm_{87.33}$
PreDyGAE $1510.41\pm_{0.01}$ $403.02\pm_{0.06}$ $529.39\pm_{58.83}$ $388.54\pm_{43.46}$ $159.45\pm_{15.71}$ $36.48\pm_{2.19}$ $162.0\pm_{20.48}$
Transformer $2154.53\pm_{30.61}$ $562.07\pm_{12.46}$ $717.62\pm_{125.1}$ $474.06\pm_{63.8}$ $179.2\pm_{25.64}$ $61.55\pm_{5.56}$ $191.91\pm_{36.18}$
Autoformer $2566.07\pm_{102.19}$ $489.69\pm_{10.56}$ $621.06\pm_{80.33}$ $387.57\pm_{33.41}$ $155.94\pm_{0.51}$ $58.1\pm_{4.51}$ $187.41\pm_{19.68}$
Informer $2212.83\pm_{2.36}$ $559.86\pm_{11.51}$ $715.12\pm_{132.16}$ $467.05\pm_{58.69}$ $178.41\pm_{28.75}$ $57.53\pm_{7.65}$ $191.85\pm_{36.5}$
Reformer $2248.94\pm_{15.68}$ $558.75\pm_{0.8}$ $713.91\pm_{131.53}$ $478.6\pm_{70.65}$ $179.85\pm_{27.33}$ $59.48\pm_{9.48}$ $195.6\pm_{31.87}$
FEDformer $2551.16\pm_{77.11}$ $501.84\pm_{4.46}$ $621.78\pm_{81.74}$ $363.84\pm_{35.37}$ $149.67\pm_{19.37}$ $57.89\pm_{2.21}$ $166.63\pm_{10.36}$
NStransformer $1337.16\pm_{47.86}$ $373.7\pm_{20.75}$ $473.97\pm_{10.71}$ $334.17\pm_{14.36}$ $128.79\pm_{8.11}$ $39.55\pm_{1.82}$ $176.86\pm_{17.5}$
PatchTST $1328.54\pm_{49.25}$ $347.07\pm_{11.24}$ $499.82\pm_{22.78}$ $321.16\pm_{15.89}$ $125.31\pm_{13.05}$ $39.74\pm_{1.04}$ $144.13\pm_{14.98}$
DLinear $1341.55\pm_{32.44}$ $350.8\pm_{6.91}$ $506.06\pm_{18.99}$ $325.15\pm_{16.25}$ $130.09\pm_{10.18}$ $38.99\pm_{1.16}$ $144.53\pm_{18.45}$
TSMixer $3040.93\pm_{86.91}$ $678.15\pm_{105.71}$ $816.57\pm_{184.07}$ $582.2\pm_{93.85}$ $211.98\pm_{22.73}$ $115.88\pm_{67.93}$ $436.64\pm_{301.19}$
FreTS $2105.11\pm_{31.62}$ $526.0\pm_{6.74}$ $679.7\pm_{130.16}$ $451.97\pm_{53.49}$ $173.34\pm_{22.83}$ $57.76\pm_{4.47}$ $188.29\pm_{25.82}$
FiLM $1298.06\pm_{265.59}$ $341.89\pm_{60.42}$ $426.14\pm_{19.46}$ $276.0\pm_{13.7}$ $111.46\pm_{9.34}$ $34.32\pm_{1.86}$ $128.54\pm_{16.33}$
Koopa $1386.85\pm_{35.21}$ $360.64\pm_{11.92}$ $532.64\pm_{14.5}$ $338.26\pm_{12.88}$ $138.63\pm_{9.35}$ $41.39\pm_{0.09}$ $152.48\pm_{17.55}$
SMAPE(%) LSTM $37.23\pm_{0.05}$ $45.67\pm_{0.02}$ $45.95\pm_{1.99}$ $50.34\pm_{1.28}$ $47.36\pm_{0.66}$ $42.47\pm_{0.4}$ $70.15\pm_{1.62}$
SegRNN $32.88\pm_{0.2}$ $43.35\pm_{0.39}$ $41.54\pm_{2.12}$ $46.46\pm_{3.08}$ $47.11\pm_{1.77}$ $44.29\pm_{4.45}$ $59.48\pm_{1.85}$
CHGH $37.21\pm_{0.01}$ $45.55\pm_{0.03}$ $46.29\pm_{1.26}$ $50.3\pm_{1.28}$ $46.22\pm_{0.72}$ $45.92\pm_{21.79}$ $77.34\pm_{37.44}$
PreDyGAE $32.0\pm_{0.02}$ $43.25\pm_{0.18}$ $49.33\pm_{2.76}$ $51.02\pm_{5.06}$ $50.59\pm_{1.35}$ $63.2\pm_{17.41}$ $70.72\pm_{0.42}$
Transformer $44.18\pm_{0.71}$ $53.31\pm_{0.19}$ $53.13\pm_{2.85}$ $56.88\pm_{3.45}$ $55.61\pm_{8.32}$ $55.47\pm_{7.12}$ $73.21\pm_{7.57}$
Autoformer $67.46\pm_{0.27}$ $68.4\pm_{0.03}$ $70.67\pm_{6.58}$ $71.92\pm_{3.02}$ $83.21\pm_{4.31}$ $83.71\pm_{4.79}$ $99.62\pm_{14.04}$
Informer $43.4\pm_{0.31}$ $53.28\pm_{0.16}$ $53.54\pm_{2.22}$ $56.42\pm_{4.14}$ $56.23\pm_{7.69}$ $54.33\pm_{8.61}$ $73.34\pm_{7.52}$
Reformer $43.1\pm_{0.01}$ $53.09\pm_{0.1}$ $53.42\pm_{2.4}$ $56.23\pm_{4.31}$ $55.84\pm_{8.2}$ $52.75\pm_{10.9}$ $76.27\pm_{3.57}$
FEDformer $68.1\pm_{0.29}$ $67.4\pm_{0.48}$ $70.57\pm_{7.06}$ $70.35\pm_{5.46}$ $76.14\pm_{3.56}$ $79.91\pm_{0.11}$ $95.19\pm_{7.37}$
NStransformer $32.85\pm_{0.14}$ $42.35\pm_{0.05}$ $44.17\pm_{1.08}$ $47.32\pm_{0.29}$ $44.23\pm_{1.08}$ $38.15\pm_{1.31}$ $103.89\pm_{60.87}$
PatchTST $31.96\pm_{2.0}$ $41.14\pm_{2.29}$ $42.63\pm_{1.59}$ $45.79\pm_{2.31}$ $43.83\pm_{2.11}$ $37.55\pm_{2.57}$ $58.04\pm_{2.54}$
DLinear $31.38\pm_{1.08}$ $40.56\pm_{1.37}$ $41.8\pm_{1.23}$ $44.83\pm_{0.38}$ $45.42\pm_{0.14}$ $41.61\pm_{0.42}$ $58.22\pm_{0.52}$
TSMixer $68.55\pm_{1.22}$ $67.06\pm_{11.47}$ $61.4\pm_{6.47}$ $75.25\pm_{0.75}$ $72.63\pm_{11.4}$ $83.29\pm_{31.13}$ $109.31\pm_{46.77}$
FreTS $37.47\pm_{0.7}$ $46.12\pm_{0.91}$ $45.36\pm_{2.66}$ $50.03\pm_{1.63}$ $46.82\pm_{1.3}$ $41.25\pm_{1.28}$ $70.0\pm_{1.47}$
FiLM $32.01\pm_{0.16}$ $41.26\pm_{0.46}$ $41.12\pm_{1.97}$ $45.56\pm_{1.21}$ $45.99\pm_{0.95}$ $40.91\pm_{0.79}$ $58.02\pm_{1.51}$
Koopa $31.01\pm_{0.02}$ $39.92\pm_{0.13}$ $42.95\pm_{3.2}$ $45.33\pm_{1.65}$ $41.88\pm_{0.65}$ $35.91\pm_{0.3}$ $57.19\pm_{1.63}$
RRMSE(%) LSTM $53.94\pm_{0.16}$ $53.9\pm_{0.12}$ $73.77\pm_{14.35}$ $98.43\pm_{12.75}$ $93.46\pm_{13.83}$ $65.39\pm_{6.3}$ $227.07\pm_{36.31}$
SegRNN $21.23\pm_{0.03}$ $22.98\pm_{0.43}$ $32.33\pm_{4.99}$ $41.84\pm_{8.49}$ $44.18\pm_{9.25}$ $34.49\pm_{6.31}$ $73.51\pm_{14.73}$
CHGH $54.21\pm_{0.03}$ $54.5\pm_{0.02}$ $77.63\pm_{10.33}$ $99.15\pm_{13.03}$ $93.96\pm_{14.27}$ $53.56\pm_{32.26}$ $105.96\pm_{104.02}$
PreDyGAE $30.72\pm_{0.01}$ $33.15\pm_{0.01}$ $46.59\pm_{9.64}$ $55.32\pm_{12.77}$ $57.56\pm_{12.47}$ $45.53\pm_{6.19}$ $91.47\pm_{23.2}$
Transformer $52.69\pm_{0.67}$ $56.3\pm_{1.62}$ $76.3\pm_{10.8}$ $99.08\pm_{12.9}$ $94.57\pm_{7.31}$ $67.56\pm_{5.78}$ $215.96\pm_{35.18}$
Autoformer $51.07\pm_{0.19}$ $39.44\pm_{1.34}$ $54.75\pm_{0.61}$ $65.4\pm_{1.15}$ $64.75\pm_{3.85}$ $54.46\pm_{8.57}$ $106.51\pm_{5.39}$
Informer $55.42\pm_{0.04}$ $56.22\pm_{2.17}$ $71.7\pm_{12.14}$ $98.81\pm_{8.31}$ $93.29\pm_{11.35}$ $64.2\pm_{6.07}$ $213.07\pm_{32.78}$
Reformer $56.91\pm_{0.64}$ $55.93\pm_{0.41}$ $76.53\pm_{13.64}$ $99.47\pm_{13.64}$ $95.12\pm_{10.02}$ $67.05\pm_{7.72}$ $221.45\pm_{36.6}$
FEDformer $49.35\pm_{2.5}$ $40.99\pm_{1.28}$ $56.8\pm_{2.22}$ $61.57\pm_{0.66}$ $63.27\pm_{4.31}$ $54.5\pm_{0.19}$ $105.93\pm_{4.82}$
NStransformer $27.47\pm_{1.23}$ $31.55\pm_{2.22}$ $43.9\pm_{3.25}$ $60.16\pm_{1.54}$ $56.5\pm_{1.0}$ $39.75\pm_{0.02}$ $116.26\pm_{2.18}$
PatchTST $27.53\pm_{1.1}$ $29.31\pm_{1.0}$ $47.82\pm_{3.34}$ $57.57\pm_{1.93}$ $55.03\pm_{2.9}$ $40.96\pm_{1.97}$ $114.76\pm_{7.12}$
DLinear $27.83\pm_{1.38}$ $29.66\pm_{1.38}$ $49.37\pm_{3.5}$ $59.59\pm_{1.4}$ $59.48\pm_{0.62}$ $40.74\pm_{0.89}$ $116.83\pm_{1.45}$
TSMixer $76.26\pm_{6.24}$ $69.83\pm_{15.63}$ $86.8\pm_{23.27}$ $119.17\pm_{17.64}$ $113.43\pm_{16.08}$ $77.61\pm_{14.65}$ $141.79\pm_{36.12}$
FreTS $52.23\pm_{1.04}$ $52.51\pm_{0.92}$ $69.87\pm_{13.95}$ $93.32\pm_{11.09}$ $88.52\pm_{11.95}$ $63.64\pm_{5.3}$ $215.64\pm_{25.29}$
FiLM $26.91\pm_{7.65}$ $28.84\pm_{7.48}$ $36.38\pm_{1.32}$ $43.93\pm_{0.54}$ $44.44\pm_{1.26}$ $31.45\pm_{0.6}$ $83.15\pm_{2.17}$
Koopa $28.98\pm_{1.27}$ $30.77\pm_{1.6}$ $53.08\pm_{5.03}$ $63.24\pm_{2.77}$ $65.24\pm_{0.87}$ $44.14\pm_{3.04}$ $129.36\pm_{2.2}$
Table S7 presents the results of repeated experiments on forecasting skill demand sequences that have undergone structural breaks. Initially, the overall errors are quite pronounced, underscoring the challenge of accurately predicting these skills. Moreover, FiLM performs well on most metrics, which further verifies the robustness of FiLM.
B.2 Job Skill Demand Forecasting for Low-Frequency Skills
In multigranular skill demand sequences, a significant number of skills remain inactive or in low frequency over extended periods. These skills might continue to have low demand in the future (indicating low importance), or they might suddenly gain interest from certain professions or companies, leading to rapid growth. In this study, we define low-frequency skills as those that appear fewer than twice in the time slices of the training set. Predicting the demand for these skills is challenging because their data points are predominantly zero during training, resulting in a lack of effective observational data. Therefore, we specifically present the demand prediction results of the existing benchmark models for these low-frequency skills, which could serve as a potential direction for further research.
Results.
We continued to test the demand prediction effect on low-frequency skills using the benchmark models described in the main text, and the results are shown in Table S8. From this, we can draw the following conclusions: Firstly, there is a significant increase in the error on the RRMSE metric, indicating that low-demand skills are difficult to predict accurately. Secondly, Koopa has the best predictive performance in this scenario. We also found that the performance of SegRNN significantly decreases, suggesting that SegRNN’s segment learning approach is not suitable for predicting low-frequency skill demands due to a lack of effective observational data, rendering the learning segments meaningless.
Table S8: Performance comparisons on skill demand series with low-frequency.
Model Market Region L1-O L2-O R&L1-O R&L2-O Company
MAE LSTM $33.69\pm_{0.44}$ $26.88\pm_{0.05}$ $16.49\pm_{0.01}$ $12.11\pm_{0.04}$ $12.28\pm_{0.02}$ $8.92\pm_{0.1}$ $12.6\pm_{0.02}$
SegRNN $51.5\pm_{0.57}$ $46.12\pm_{0.28}$ $25.65\pm_{0.39}$ $18.03\pm_{0.03}$ $22.08\pm_{1.92}$ $19.1\pm_{3.22}$ $19.37\pm_{0.88}$
CHGH $32.37\pm_{0.0}$ $24.99\pm_{0.02}$ $14.2\pm_{0.04}$ $9.32\pm_{0.01}$ $9.45\pm_{0.0}$ $67.89\pm_{10.03}$ $77.01\pm_{5.58}$
PreDyGAE $113.67\pm_{1.73}$ $157.17\pm_{18.68}$ $343.32\pm_{0.15}$ $64.46\pm_{13.49}$ $24.16\pm_{0.13}$ $60.41\pm_{55.15}$ $68.89\pm_{62.89}$
Transformer $54.84\pm_{1.99}$ $52.95\pm_{0.09}$ $46.1\pm_{0.13}$ $45.47\pm_{0.07}$ $45.64\pm_{0.02}$ $45.1\pm_{0.01}$ $45.43\pm_{0.0}$
Autoformer $132.98\pm_{11.42}$ $106.13\pm_{1.65}$ $110.06\pm_{0.19}$ $97.66\pm_{1.23}$ $98.28\pm_{0.46}$ $95.26\pm_{1.1}$ $90.23\pm_{0.54}$
Informer $54.42\pm_{1.83}$ $52.76\pm_{0.13}$ $46.1\pm_{0.0}$ $45.52\pm_{0.01}$ $45.69\pm_{0.02}$ $45.07\pm_{0.01}$ $45.47\pm_{0.0}$
Reformer $52.72\pm_{0.3}$ $53.25\pm_{0.15}$ $46.19\pm_{0.07}$ $45.58\pm_{0.02}$ $45.65\pm_{0.01}$ $45.08\pm_{0.03}$ $45.49\pm_{0.03}$
FEDformer $133.07\pm_{6.09}$ $101.47\pm_{1.47}$ $109.51\pm_{1.3}$ $96.43\pm_{1.0}$ $95.83\pm_{1.48}$ $94.18\pm_{0.48}$ $90.71\pm_{0.47}$
NStransformer $57.96\pm_{1.83}$ $34.24\pm_{0.3}$ $17.99\pm_{0.37}$ $10.84\pm_{0.04}$ $11.46\pm_{0.02}$ $6.45\pm_{0.02}$ $11.9\pm_{0.0}$
PatchTST $47.25\pm_{6.07}$ $33.21\pm_{3.3}$ $17.31\pm_{1.85}$ $10.39\pm_{1.03}$ $10.91\pm_{1.1}$ $6.1\pm_{0.67}$ $11.4\pm_{1.23}$
DLinear $51.7\pm_{7.31}$ $42.08\pm_{8.83}$ $29.64\pm_{9.87}$ $25.18\pm_{10.9}$ $25.6\pm_{10.9}$ $22.53\pm_{11.72}$ $25.79\pm_{10.77}$
TSMixer $102.23\pm_{14.26}$ $64.77\pm_{10.8}$ $51.89\pm_{9.94}$ $52.61\pm_{1.14}$ $40.06\pm_{7.0}$ $34.02\pm_{2.2}$ $41.32\pm_{2.03}$
FreTS $38.55\pm_{0.83}$ $28.89\pm_{0.96}$ $16.6\pm_{0.58}$ $11.19\pm_{0.47}$ $11.46\pm_{0.51}$ $7.44\pm_{0.47}$ $11.94\pm_{0.48}$
FiLM $51.42\pm_{20.47}$ $34.97\pm_{10.99}$ $18.23\pm_{5.69}$ $10.9\pm_{3.11}$ $11.47\pm_{3.38}$ $6.43\pm_{1.96}$ $11.95\pm_{3.58}$
Koopa $43.3\pm_{4.56}$ $29.66\pm_{2.37}$ $15.52\pm_{1.38}$ $9.49\pm_{0.73}$ $9.95\pm_{0.82}$ $5.53\pm_{0.48}$ $10.35\pm_{0.86}$
RMSE LSTM $99.83\pm_{0.77}$ $306.02\pm_{0.33}$ $168.19\pm_{0.01}$ $140.57\pm_{0.0}$ $247.24\pm_{0.02}$ $125.6\pm_{0.01}$ $272.77\pm_{0.04}$
SegRNN $141.21\pm_{0.31}$ $452.49\pm_{0.09}$ $233.37\pm_{0.06}$ $146.34\pm_{0.22}$ $216.77\pm_{0.01}$ $117.81\pm_{0.63}$ $227.59\pm_{0.0}$
CHGH $100.44\pm_{0.02}$ $306.28\pm_{0.01}$ $168.23\pm_{0.01}$ $140.54\pm_{0.0}$ $247.21\pm_{0.01}$ $134.20\pm_{0.0}$ $208.01\pm_{17.09}$
PreDyGAE $169.76\pm_{1.46}$ $542.42\pm_{24.84}$ $771.72\pm_{50.15}$ $195.02\pm_{16.23}$ $218.94\pm_{0.65}$ $119.76\pm_{109.33}$ $188.17\pm_{171.77}$
Transformer $112.3\pm_{5.16}$ $311.19\pm_{1.43}$ $176.74\pm_{0.52}$ $150.63\pm_{0.53}$ $252.9\pm_{0.1}$ $137.39\pm_{0.13}$ $278.45\pm_{0.13}$
Autoformer $249.03\pm_{36.15}$ $364.32\pm_{9.52}$ $231.67\pm_{3.95}$ $184.29\pm_{1.11}$ $266.1\pm_{0.64}$ $172.38\pm_{1.84}$ $281.53\pm_{0.17}$
Informer $111.59\pm_{6.84}$ $312.15\pm_{0.35}$ $176.18\pm_{0.19}$ $150.46\pm_{0.16}$ $252.84\pm_{0.01}$ $137.07\pm_{0.12}$ $279.56\pm_{0.36}$
Reformer $107.11\pm_{3.02}$ $314.04\pm_{1.77}$ $176.73\pm_{0.57}$ $150.62\pm_{0.23}$ $252.85\pm_{0.14}$ $137.02\pm_{0.12}$ $279.79\pm_{0.61}$
FEDformer $233.86\pm_{32.3}$ $346.25\pm_{1.54}$ $227.64\pm_{2.23}$ $182.54\pm_{0.06}$ $264.12\pm_{0.64}$ $171.43\pm_{0.76}$ $280.57\pm_{1.13}$
NStransformer $159.4\pm_{10.66}$ $329.86\pm_{4.31}$ $182.83\pm_{8.95}$ $136.01\pm_{2.69}$ $249.63\pm_{0.44}$ $129.74\pm_{5.67}$ $255.48\pm_{4.76}$
PatchTST $126.71\pm_{17.24}$ $339.92\pm_{16.66}$ $182.49\pm_{8.84}$ $134.57\pm_{2.89}$ $233.14\pm_{1.45}$ $120.94\pm_{1.7}$ $253.85\pm_{3.35}$
DLinear $109.06\pm_{5.5}$ $320.09\pm_{2.51}$ $173.51\pm_{2.67}$ $133.82\pm_{2.37}$ $235.23\pm_{1.35}$ $123.01\pm_{2.84}$ $254.7\pm_{1.91}$
TSMixer $186.01\pm_{18.76}$ $359.57\pm_{28.14}$ $199.07\pm_{10.44}$ $160.52\pm_{5.4}$ $261.65\pm_{6.67}$ $139.7\pm_{3.15}$ $296.15\pm_{1.83}$
FreTS $104.43\pm_{1.38}$ $310.94\pm_{2.43}$ $170.36\pm_{1.47}$ $139.96\pm_{0.74}$ $239.77\pm_{4.49}$ $122.58\pm_{2.07}$ $266.82\pm_{3.88}$
FiLM $129.11\pm_{32.28}$ $343.63\pm_{34.15}$ $185.3\pm_{16.68}$ $135.8\pm_{1.72}$ $235.13\pm_{10.73}$ $122.84\pm_{2.06}$ $256.54\pm_{7.91}$
Koopa $111.96\pm_{8.52}$ $317.8\pm_{3.82}$ $171.28\pm_{3.23}$ $134.11\pm_{0.67}$ $242.79\pm_{2.91}$ $124.64\pm_{0.36}$ $260.87\pm_{0.46}$
SMAPE(%) LSTM $26.1\pm_{0.39}$ $19.43\pm_{0.17}$ $15.18\pm_{0.01}$ $12.88\pm_{0.07}$ $13.01\pm_{0.02}$ $11.46\pm_{0.17}$ $13.01\pm_{0.03}$
SegRNN $33.71\pm_{0.71}$ $29.55\pm_{0.4}$ $22.04\pm_{0.56}$ $18.44\pm_{0.11}$ $24.16\pm_{2.97}$ $24.8\pm_{4.63}$ $19.83\pm_{1.32}$
CHGH $23.61\pm_{0.0}$ $15.99\pm_{0.04}$ $11.08\pm_{0.09}$ $7.8\pm_{0.03}$ $7.86\pm_{0.01}$ $9.61\pm_{1.08}$ $8.07\pm_{2.03}$
PreDyGAE $83.59\pm_{0.72}$ $82.33\pm_{4.38}$ $114.25\pm_{1.42}$ $48.0\pm_{7.33}$ $23.97\pm_{0.1}$ $47.34\pm_{43.22}$ $48.31\pm_{44.1}$
Transformer $49.05\pm_{1.26}$ $48.23\pm_{0.0}$ $48.68\pm_{0.14}$ $50.78\pm_{0.03}$ $50.92\pm_{0.01}$ $52.6\pm_{0.01}$ $50.28\pm_{0.0}$
Autoformer $80.95\pm_{1.5}$ $76.73\pm_{0.44}$ $81.69\pm_{0.03}$ $80.14\pm_{0.55}$ $80.43\pm_{0.21}$ $81.25\pm_{0.49}$ $76.73\pm_{0.25}$
Informer $48.87\pm_{0.78}$ $48.07\pm_{0.25}$ $48.69\pm_{0.03}$ $50.83\pm_{0.01}$ $50.96\pm_{0.02}$ $52.57\pm_{0.01}$ $50.29\pm_{0.0}$
Reformer $48.32\pm_{1.06}$ $48.33\pm_{0.02}$ $48.76\pm_{0.08}$ $50.87\pm_{0.03}$ $50.93\pm_{0.0}$ $52.59\pm_{0.02}$ $50.28\pm_{0.02}$
FEDformer $82.56\pm_{1.23}$ $75.32\pm_{0.52}$ $81.88\pm_{0.57}$ $79.7\pm_{0.39}$ $79.55\pm_{0.62}$ $80.83\pm_{0.21}$ $76.96\pm_{0.21}$
NStransformer $32.62\pm_{0.69}$ $18.84\pm_{0.31}$ $11.59\pm_{0.02}$ $7.3\pm_{0.01}$ $7.81\pm_{0.01}$ $4.76\pm_{0.0}$ $7.76\pm_{0.0}$
PatchTST $29.48\pm_{1.53}$ $18.23\pm_{1.46}$ $11.15\pm_{1.04}$ $7.01\pm_{0.71}$ $7.51\pm_{0.76}$ $4.56\pm_{0.52}$ $7.46\pm_{0.8}$
DLinear $42.19\pm_{8.16}$ $35.1\pm_{10.8}$ $31.64\pm_{12.55}$ $30.55\pm_{14.04}$ $30.93\pm_{14.03}$ $30.35\pm_{15.14}$ $30.52\pm_{13.82}$
TSMixer $68.79\pm_{5.48}$ $51.84\pm_{6.45}$ $48.47\pm_{6.43}$ $54.61\pm_{0.92}$ $45.87\pm_{6.93}$ $43.34\pm_{2.84}$ $45.68\pm_{1.95}$
FreTS $29.63\pm_{0.33}$ $19.71\pm_{0.62}$ $14.01\pm_{0.56}$ $10.6\pm_{0.64}$ $10.86\pm_{0.67}$ $8.45\pm_{0.75}$ $11.09\pm_{0.61}$
FiLM $30.74\pm_{10.37}$ $18.56\pm_{6.37}$ $11.42\pm_{4.06}$ $7.19\pm_{2.55}$ $7.7\pm_{2.79}$ $4.69\pm_{1.69}$ $7.64\pm_{2.84}$
Koopa $28.27\pm_{2.34}$ $16.67\pm_{1.38}$ $10.15\pm_{0.9}$ $6.35\pm_{0.57}$ $6.79\pm_{0.63}$ $4.09\pm_{0.39}$ $6.73\pm_{0.63}$
RRMSE(%) LSTM $432.63\pm_{21.64}$ $1364.7\pm_{202.93}$ $1057.49\pm_{28.76}$ $1468.93\pm_{15.86}$ $2581.13\pm_{33.65}$ $1485.07\pm_{28.13}$ $1828.92\pm_{33.42}$
SegRNN $114.7\pm_{0.21}$ $125.97\pm_{0.16}$ $126.24\pm_{0.18}$ $146.81\pm_{0.73}$ $130.23\pm_{0.45}$ $130.24\pm_{0.87}$ $133.03\pm_{0.54}$
CHGH $452.99\pm_{0.65}$ $1662.7\pm_{2.11}$ $1394.35\pm_{4.28}$ $1824.0\pm_{1.1}$ $3436.23\pm_{14.46}$ $289.09\pm_{14.07}$ $406.07\pm_{8.08}$
PreDyGAE $104.17\pm_{0.99}$ $216.23\pm_{2.0}$ $201.73\pm_{0.3}$ $220.44\pm_{4.62}$ $229.42\pm_{0.22}$ $165.21\pm_{59.52}$ $272.3\pm_{66.0}$
Transformer $172.07\pm_{30.78}$ $513.13\pm_{29.31}$ $320.11\pm_{12.98}$ $269.97\pm_{3.45}$ $468.03\pm_{0.41}$ $242.99\pm_{4.1}$ $466.5\pm_{1.48}$
Autoformer $89.47\pm_{0.4}$ $161.16\pm_{10.32}$ $136.15\pm_{6.04}$ $134.91\pm_{1.17}$ $195.47\pm_{1.88}$ $134.43\pm_{1.31}$ $191.17\pm_{2.0}$
Informer $193.88\pm_{30.22}$ $538.02\pm_{31.16}$ $316.88\pm_{4.27}$ $275.64\pm_{0.03}$ $470.15\pm_{2.27}$ $246.67\pm_{0.09}$ $462.58\pm_{1.78}$
Reformer $200.52\pm_{32.93}$ $536.21\pm_{13.15}$ $324.86\pm_{11.04}$ $276.66\pm_{0.14}$ $470.18\pm_{0.3}$ $247.32\pm_{0.89}$ $460.86\pm_{4.08}$
FEDformer $96.72\pm_{6.76}$ $176.64\pm_{0.39}$ $136.62\pm_{1.69}$ $136.93\pm_{0.31}$ $199.25\pm_{1.04}$ $134.8\pm_{0.73}$ $189.29\pm_{1.49}$
NStransformer $90.14\pm_{7.16}$ $219.18\pm_{39.19}$ $193.48\pm_{36.91}$ $218.73\pm_{0.68}$ $406.29\pm_{7.04}$ $272.29\pm_{31.63}$ $259.15\pm_{8.34}$
PatchTST $95.85\pm_{5.0}$ $210.2\pm_{36.63}$ $199.12\pm_{34.35}$ $266.55\pm_{38.3}$ $391.36\pm_{46.33}$ $298.48\pm_{41.01}$ $278.87\pm_{33.7}$
DLinear $90.72\pm_{2.96}$ $241.41\pm_{14.46}$ $219.68\pm_{17.61}$ $281.05\pm_{37.87}$ $431.86\pm_{48.87}$ $296.1\pm_{49.89}$ $300.39\pm_{25.41}$
TSMixer $125.99\pm_{9.23}$ $200.39\pm_{22.34}$ $211.03\pm_{36.38}$ $192.85\pm_{35.49}$ $418.03\pm_{16.7}$ $275.29\pm_{1.9}$ $293.38\pm_{24.47}$
FreTS $222.55\pm_{30.4}$ $479.45\pm_{20.87}$ $462.19\pm_{33.86}$ $720.43\pm_{51.2}$ $962.73\pm_{47.52}$ $758.24\pm_{41.67}$ $809.76\pm_{76.22}$
FiLM $108.05\pm_{34.05}$ $444.72\pm_{401.85}$ $360.89\pm_{298.05}$ $402.84\pm_{297.56}$ $652.7\pm_{546.99}$ $417.27\pm_{297.85}$ $364.47\pm_{235.25}$
Koopa $93.83\pm_{0.11}$ $287.42\pm_{57.85}$ $264.45\pm_{56.56}$ $345.86\pm_{79.42}$ $582.52\pm_{154.25}$ $368.66\pm_{64.94}$ $345.87\pm_{66.81}$
B.3 Skill Co-occurrence Graph Enhanced Job Skill Demand Forecasting
In the task of job skill demand forecasting, fully leveraging the inter-relationships among different skills is beneficial for downstream tasks. Therefore, we construct a prior graph with co-occurrence frequency from the training data to include as a dataset component. Given a set of granularities $i,j,\ldots,k$, we constructed the skill co-occurrence graph as $\mathcal{G}^{i,j,\ldots,k}=(\mathcal{V}^{i,j,\ldots,k},\mathcal{E}^{i,j,\ldots,k})$, where $\mathcal{V}^{i,j,\ldots,k}$ is the extended skill set under the multiple granularities. The edge weight $e_{v,v^{\prime}}\in\mathcal{E}^{i,j,...,k}$ between nodes $v$ and $v^{\prime}$ is determined by the co-occurrence frequency of the node pair $v,v^{\prime}$ in the job advertisement data for training $\mathcal{P}^{train}$. Specifically, given $v=(a^{i},a^{j},\ldots,a^{k},s)$ and $v^{\prime}=(a^{i\prime},a^{j\prime},...,a^{k\prime},s^{\prime})$, $e_{v,v^{\prime}}$ is calculated as:
$e_{v,v^{\prime}}=\sum_{p\in\mathcal{P}^{train}}\prod_{x\in\{a^{i},a^{j},\ldots,a^{k},a^{i\prime},a^{j\prime},\ldots,a^{k\prime},s,s^{\prime}\}}\mathbf{1}_{p}(x\in p).$ (4)
This information will serve as prior knowledge, reflecting global inter-skill dependency patterns.
Benchmark Models
To fully utilize the prior information from the co-occurency graph, we introduce several GNN-based methods for multivariate time series prediction. These methods leverage GNNs to extract the influences between different variables, effectively capturing the relationships among various time series. The specific models are as follows:
  • EvolveGCN [42]: EvolveGCN introduces a recurrent mechanism to update the network parameters, as GCN parameters, for capturing the dynamism of the graphs. Two methods are introduced: EvolveGCNH, which learns the weight matrix of the graph at each time step as a hidden state, and EvolveGCNO, which directly employs the weight evolution as a hidden state output, decoupled from node embedding.
  • GConvGRU [43]: This model integrates convolutional neural networks (CNNs) on graphs to identify spatial structures and recurrent neural networks (RNNs) to detect dynamic patterns. Two architectures, GConvGRU and GConvLSTM, are explored for the Graph Convolutional Recurrent Network (GCRN).
  • TGCN [44]: The temporal graph convolutional network (T-GCN) model, which is in combination with the graph convolutional network (GCN) and gated recurrent unit (GRU). Specifically, the GCN is used to learn complex topological structures to capture spatial dependence and the gated recurrent unit is used to learn dynamic changes of traffic data to capture temporal dependence.
  • GCLSTM [45]: GCLSTM is an end-to-end model integrating a Graph Convolution Network (GCN) embedded Long Short-Term Memory network (LSTM) for dynamic network link prediction. The GCN captures local structural properties, while the LSTM learns temporal features across snapshots of a dynamic network.
  • DyGrEncoder [46]: This approach combines a sequence-to-sequence encoder-decoder model with gated graph neural networks (GGNNs) and long short-term memory networks (LSTMs). The encoder captures temporal dynamics in an evolving graph, and the decoder reconstructs the dynamics using the encoded representation.
We implement these benchmark models using the PyG library  [^3] https://github.com/benedekrozemberczki/pytorch_geometric_temporal and demonstrate the effectiveness of these GNN-based methods in skill demand forecasting.
Results
We have implemented a series of graph-based multivariate time series forecasting methods based on the co-occurrence graph and verified their experimental effects under the three scenarios discussed above. Firstly, Table S9 presents the overall performance of the methods based on the co-occurrence graph for skill demand forecasting. It is observed that the prediction accuracy significantly declines across the overall labor market. However, as the granularity of the forecast becomes finer, the model performance improves, and at a finer granularity, the EvolveGCN method outperforms the state-of-the-art (SOTA) methods mentioned in the main text considerably. We analyze that the finer the granularity, the more accurately the co-occurrence graph reflects the associations between skills, while coarser granularity might introduce excessive noise leading to decreased model performance. The fine-grained co-occurrence graph accurately reflects the interrelationships between skills at different granularities, which aids in enhancing the model’s prediction accuracy. Secondly, we find significant differences in the RRMSE metric among these methods, with EvolveGCN showing superior performance because it can learn the evolution of GCN parameter weights over time, thus capturing the evolving dependencies among edges. Therefore, based on the provided co-occurrence graph, it can effectively learn the evolution of skill relationships, which is beneficial for dynamic prediction of skill demand.
For skill demand forecasting in scenarios involving structural breaks, as shown in Table S10, the improvements in the methods based on the co-occurrence graph are greater than those in the overall skill demand forecasting task. This suggests that skills experiencing structural breaks have strong interconnections, and the co-occurrence graph helps the model to identify the patterns of skill demand sequences that are likely to undergo structural breaks, thus further enhancing the prediction effectiveness for this category of skills.
In the task of predicting low-frequency skills, as shown in Table S11, methods like GConvLSTM significantly outperform EvolveGCN. This is due to the sparse observable data for these skills, which leads to sparse connectivity edges on the co-occurrence graph.
Table S9: Performance comparisons on skill demand series with GNN-based methods.
Model Market Region L1-O L2-O R&L1-O R&L2-O Company
MAE EvolveGCNH $1053.18_{\pm 804.38}$ $30.07_{\pm 22.97}$ $16.51_{\pm 12.61}$ $5.43_{\pm 4.15}$ $2.92_{\pm 2.23}$ $1.11_{\pm 0.85}$ $0.96_{\pm 0.73}$
EvolveGCNO $151.13_{\pm 115.43}$ $27.01_{\pm 20.63}$ $15.4_{\pm 11.76}$ $5.04_{\pm 3.85}$ $2.95_{\pm 2.25}$ $1.21_{\pm 0.92}$ $0.89_{\pm 0.68}$
GConvGRU $570.31_{\pm 435.58}$ $92.96_{\pm 71.0}$ $46.43_{\pm 35.46}$ $11.45_{\pm 8.75}$ $5.47_{\pm 4.18}$ $1.4_{\pm 1.07}$ $1.04_{\pm 0.79}$
TGCN $741.05_{\pm 565.99}$ $96.43_{\pm 73.65}$ $55.41_{\pm 42.32}$ $13.21_{\pm 10.09}$ $6.28_{\pm 4.8}$ $1.63_{\pm 1.24}$ $1.11_{\pm 0.85}$
GCLSTM $729.48_{\pm 557.15}$ $57.13_{\pm 43.63}$ $25.84_{\pm 19.74}$ $11.49_{\pm 8.78}$ $5.67_{\pm 4.33}$ $1.61_{\pm 1.23}$ $1.07_{\pm 0.82}$
GConvLSTM $741.45_{\pm 566.29}$ $93.61_{\pm 71.5}$ $46.57_{\pm 35.57}$ $11.19_{\pm 8.55}$ $5.57_{\pm 4.25}$ $1.49_{\pm 1.14}$ $1.09_{\pm 0.83}$
DyGrEncoder $732.53_{\pm 559.48}$ $92.21_{\pm 70.43}$ $46.96_{\pm 35.87}$ $12.26_{\pm 9.36}$ $5.57_{\pm 4.26}$ $1.47_{\pm 1.13}$ $1.07_{\pm 0.82}$
RMSE EvolveGCNH $4998.7_{\pm 3817.82}$ $166.34_{\pm 127.04}$ $178.55_{\pm 136.37}$ $76.14_{\pm 58.15}$ $35.12_{\pm 26.82}$ $16.97_{\pm 12.96}$ $19.33_{\pm 14.77}$
EvolveGCNO $709.19_{\pm 541.65}$ $143.72_{\pm 109.77}$ $164.16_{\pm 125.38}$ $71.57_{\pm 54.67}$ $33.21_{\pm 25.37}$ $15.04_{\pm 11.49}$ $17.93_{\pm 13.7}$
GConvGRU $2968.75_{\pm 2267.42}$ $579.66_{\pm 442.72}$ $442.53_{\pm 337.99}$ $170.09_{\pm 129.91}$ $81.95_{\pm 62.59}$ $31.75_{\pm 24.25}$ $30.65_{\pm 23.41}$
TGCN $3163.46_{\pm 2416.13}$ $581.49_{\pm 444.12}$ $450.79_{\pm 344.29}$ $172.13_{\pm 131.47}$ $83.33_{\pm 63.64}$ $32.19_{\pm 24.59}$ $30.77_{\pm 23.5}$
GCLSTM $3159.61_{\pm 2413.19}$ $497.97_{\pm 380.33}$ $371.13_{\pm 283.45}$ $170.23_{\pm 130.01}$ $82.66_{\pm 63.13}$ $32.25_{\pm 24.63}$ $30.81_{\pm 23.53}$
GConvLSTM $3166.41_{\pm 2418.38}$ $581.26_{\pm 443.94}$ $442.72_{\pm 338.13}$ $169.43_{\pm 129.4}$ $82.38_{\pm 62.92}$ $32.29_{\pm 24.66}$ $30.88_{\pm 23.58}$
DyGrEncoder $3161.61_{\pm 2414.72}$ $579.78_{\pm 442.81}$ $443.49_{\pm 338.72}$ $171.91_{\pm 131.3}$ $82.27_{\pm 62.83}$ $32.02_{\pm 24.46}$ $30.76_{\pm 23.49}$
SMAPE(%) EvolveGCNH $76.69_{\pm 58.58}$ $41.67_{\pm 31.83}$ $39.77_{\pm 30.38}$ $33.11_{\pm 25.29}$ $28.99_{\pm 22.14}$ $18.79_{\pm 14.35}$ $28.36_{\pm 21.66}$
EvolveGCNO $45.32_{\pm 34.61}$ $56.75_{\pm 43.35}$ $46.34_{\pm 35.39}$ $35.09_{\pm 26.8}$ $29.33_{\pm 22.4}$ $33.19_{\pm 25.35}$ $25.99_{\pm 19.85}$
GConvGRU $63.93_{\pm 48.82}$ $62.41_{\pm 47.67}$ $38.65_{\pm 29.52}$ $31.09_{\pm 23.74}$ $23.67_{\pm 18.08}$ $12.93_{\pm 9.88}$ $20.5_{\pm 15.66}$
TGCN $71.13_{\pm 54.32}$ $51.37_{\pm 39.23}$ $56.23_{\pm 42.95}$ $37.21_{\pm 28.42}$ $25.83_{\pm 19.73}$ $21.74_{\pm 16.6}$ $23.35_{\pm 17.83}$
GCLSTM $60.49_{\pm 46.2}$ $38.68_{\pm 29.54}$ $43.6_{\pm 33.3}$ $28.15_{\pm 21.5}$ $24.34_{\pm 18.59}$ $18.57_{\pm 14.18}$ $21.74_{\pm 16.6}$
GConvLSTM $64.79_{\pm 49.49}$ $44.77_{\pm 34.2}$ $35.9_{\pm 27.42}$ $26.21_{\pm 20.02}$ $21.67_{\pm 16.55}$ $13.95_{\pm 10.66}$ $23.05_{\pm 17.61}$
DyGrEncoder $61.55_{\pm 47.01}$ $44.74_{\pm 34.17}$ $38.12_{\pm 29.11}$ $26.94_{\pm 20.58}$ $26.49_{\pm 20.23}$ $18.42_{\pm 14.07}$ $23.75_{\pm 18.14}$
RRMSE(%) EvolveGCNH $68.68_{\pm 52.46}$ $19.57_{\pm 14.95}$ $32.99_{\pm 25.2}$ $36.53_{\pm 27.9}$ $32.1_{\pm 24.52}$ $41.01_{\pm 31.32}$ $65.93_{\pm 50.35}$
EvolveGCNO $16.21_{\pm 12.38}$ $17.85_{\pm 13.63}$ $28.93_{\pm 22.09}$ $34.6_{\pm 26.43}$ $30.4_{\pm 23.22}$ $32.27_{\pm 24.64}$ $58.07_{\pm 44.35}$
GConvGRU $418.84_{\pm 319.89}$ $989.03_{\pm 755.38}$ $1111.69_{\pm 849.06}$ $701.09_{\pm 535.46}$ $414.17_{\pm 316.33}$ $307.07_{\pm 234.53}$ $385.69_{\pm 294.58}$
TGCN $3478.62_{\pm 2656.84}$ $1037.13_{\pm 792.12}$ $112164.83_{\pm 85667.3}$ $948.04_{\pm 724.08}$ $501.07_{\pm 382.7}$ $338.75_{\pm 258.73}$ $404.5_{\pm 308.94}$
GCLSTM $3083.62_{\pm 2355.15}$ $176.53_{\pm 134.83}$ $159.96_{\pm 122.17}$ $721.9_{\pm 551.36}$ $467.49_{\pm 357.05}$ $334.26_{\pm 255.3}$ $425.46_{\pm 324.95}$
GConvLSTM $4479.9_{\pm 3421.58}$ $1123.32_{\pm 857.95}$ $1101.0_{\pm 840.9}$ $640.31_{\pm 489.05}$ $449.12_{\pm 343.02}$ $382.23_{\pm 291.94}$ $451.71_{\pm 345.0}$
DyGrEncoder $3386.49_{\pm 2586.47}$ $1010.93_{\pm 772.11}$ $1247.51_{\pm 952.8}$ $1004.62_{\pm 767.29}$ $450.57_{\pm 344.13}$ $331.31_{\pm 253.04}$ $418.67_{\pm 319.77}$
Table S10: Performance comparisons on skill demand series with structural breaks with GNN-based methods.
Model Market Region L1-O L2-O R&L1-O R&L2-O Company
MAE EvolveGCNH $1385.2_{\pm 1057.96}$ $61.79_{\pm 47.19}$ $58.05_{\pm 44.33}$ $36.04_{\pm 27.53}$ $10.69_{\pm 8.17}$ $5.56_{\pm 4.25}$ $13.25_{\pm 10.12}$
EvolveGCNO $195.87_{\pm 149.6}$ $53.86_{\pm 41.14}$ $52.46_{\pm 40.07}$ $32.76_{\pm 25.02}$ $10.85_{\pm 8.28}$ $5.21_{\pm 3.98}$ $12.22_{\pm 9.33}$
GConvGRU $749.83_{\pm 572.69}$ $204.83_{\pm 156.44}$ $189.48_{\pm 144.72}$ $92.29_{\pm 70.49}$ $26.94_{\pm 20.58}$ $9.23_{\pm 7.05}$ $17.26_{\pm 13.18}$
TGCN $990.21_{\pm 756.28}$ $210.7_{\pm 160.92}$ $222.48_{\pm 169.92}$ $100.93_{\pm 77.09}$ $29.27_{\pm 22.36}$ $9.93_{\pm 7.58}$ $17.87_{\pm 13.65}$
GCLSTM $977.98_{\pm 746.94}$ $128.33_{\pm 98.02}$ $102.44_{\pm 78.24}$ $92.71_{\pm 70.81}$ $27.94_{\pm 21.34}$ $10.25_{\pm 7.83}$ $17.81_{\pm 13.61}$
GConvLSTM $993.55_{\pm 758.84}$ $207.86_{\pm 158.76}$ $190.31_{\pm 145.35}$ $90.33_{\pm 68.99}$ $27.51_{\pm 21.01}$ $9.93_{\pm 7.58}$ $18.0_{\pm 13.75}$
DyGrEncoder $982.05_{\pm 750.06}$ $204.89_{\pm 156.49}$ $191.91_{\pm 146.57}$ $98.78_{\pm 75.44}$ $27.41_{\pm 20.93}$ $9.6_{\pm 7.33}$ $17.51_{\pm 13.38}$
RMSE EvolveGCNH $5799.81_{\pm 4429.68}$ $261.83_{\pm 199.97}$ $371.99_{\pm 284.11}$ $227.88_{\pm 174.05}$ $87.98_{\pm 67.2}$ $29.87_{\pm 22.82}$ $100.77_{\pm 76.96}$
EvolveGCNO $863.11_{\pm 659.21}$ $226.98_{\pm 173.36}$ $341.34_{\pm 260.7}$ $213.99_{\pm 163.44}$ $82.25_{\pm 62.82}$ $25.03_{\pm 19.11}$ $93.36_{\pm 71.3}$
GConvGRU $3463.23_{\pm 2645.08}$ $890.59_{\pm 680.2}$ $922.56_{\pm 704.62}$ $508.61_{\pm 388.46}$ $198.01_{\pm 151.24}$ $69.96_{\pm 53.43}$ $158.49_{\pm 121.05}$
TGCN $3688.76_{\pm 2817.34}$ $893.17_{\pm 682.17}$ $938.93_{\pm 717.12}$ $514.25_{\pm 392.77}$ $201.03_{\pm 153.54}$ $71.44_{\pm 54.56}$ $159.07_{\pm 121.49}$
GCLSTM $3684.35_{\pm 2813.97}$ $768.16_{\pm 586.69}$ $776.42_{\pm 593.0}$ $508.99_{\pm 388.74}$ $199.63_{\pm 152.47}$ $71.71_{\pm 54.77}$ $159.26_{\pm 121.64}$
GConvLSTM $3692.17_{\pm 2819.94}$ $892.93_{\pm 681.98}$ $922.93_{\pm 704.9}$ $506.71_{\pm 387.0}$ $198.99_{\pm 151.98}$ $71.93_{\pm 54.94}$ $159.59_{\pm 121.89}$
DyGrEncoder $3686.65_{\pm 2815.72}$ $890.76_{\pm 680.33}$ $924.49_{\pm 706.09}$ $513.77_{\pm 392.4}$ $198.73_{\pm 151.79}$ $70.95_{\pm 54.19}$ $159.01_{\pm 121.44}$
SMAPE(%) EvolveGCNH $77.85_{\pm 59.46}$ $31.5_{\pm 24.06}$ $30.67_{\pm 23.43}$ $33.05_{\pm 25.24}$ $33.43_{\pm 25.54}$ $29.61_{\pm 22.61}$ $38.66_{\pm 29.53}$
EvolveGCNO $31.65_{\pm 24.17}$ $33.88_{\pm 25.88}$ $30.29_{\pm 23.14}$ $31.57_{\pm 24.11}$ $34.17_{\pm 26.1}$ $37.43_{\pm 28.59}$ $37.69_{\pm 28.78}$
GConvGRU $49.4_{\pm 37.73}$ $44.31_{\pm 33.84}$ $44.6_{\pm 34.06}$ $38.91_{\pm 29.72}$ $34.25_{\pm 26.16}$ $27.57_{\pm 21.06}$ $37.43_{\pm 28.59}$
TGCN $71.56_{\pm 54.65}$ $49.2_{\pm 37.58}$ $120.31_{\pm 91.89}$ $48.37_{\pm 36.94}$ $36.76_{\pm 28.08}$ $32.09_{\pm 24.51}$ $39.02_{\pm 29.8}$
GCLSTM $63.9_{\pm 48.8}$ $31.28_{\pm 23.89}$ $28.84_{\pm 22.03}$ $39.16_{\pm 29.91}$ $35.03_{\pm 26.76}$ $32.18_{\pm 24.58}$ $37.87_{\pm 28.93}$
GConvLSTM $68.73_{\pm 52.49}$ $45.39_{\pm 34.66}$ $44.43_{\pm 33.94}$ $38.47_{\pm 29.38}$ $33.75_{\pm 25.77}$ $28.62_{\pm 21.86}$ $37.96_{\pm 28.99}$
DyGrEncoder $64.22_{\pm 49.05}$ $43.72_{\pm 33.39}$ $44.73_{\pm 34.16}$ $43.73_{\pm 33.4}$ $35.73_{\pm 27.29}$ $30.42_{\pm 23.23}$ $37.41_{\pm 28.57}$
RRMSE(%) EvolveGCNH $68.88_{\pm 52.61}$ $20.26_{\pm 15.47}$ $33.11_{\pm 25.29}$ $36.88_{\pm 28.17}$ $35.03_{\pm 26.76}$ $26.91_{\pm 20.55}$ $68.99_{\pm 52.69}$
EvolveGCNO $17.01_{\pm 12.99}$ $18.47_{\pm 14.11}$ $28.97_{\pm 22.12}$ $34.88_{\pm 26.64}$ $32.23_{\pm 24.61}$ $20.17_{\pm 15.41}$ $59.82_{\pm 45.69}$
GConvGRU $426.16_{\pm 325.49}$ $1072.11_{\pm 818.84}$ $1215.75_{\pm 928.55}$ $763.66_{\pm 583.25}$ $485.64_{\pm 370.91}$ $255.02_{\pm 194.77}$ $427.8_{\pm 326.74}$
TGCN $3708.55_{\pm 2832.45}$ $1159.31_{\pm 885.44}$ $156814.93_{\pm 119769.38}$ $1119.75_{\pm 855.22}$ $621.0_{\pm 474.3}$ $290.69_{\pm 222.02}$ $451.93_{\pm 345.17}$
GCLSTM $3210.27_{\pm 2451.88}$ $184.79_{\pm 141.14}$ $164.56_{\pm 125.68}$ $786.29_{\pm 600.54}$ $553.67_{\pm 422.87}$ $291.57_{\pm 222.69}$ $475.27_{\pm 362.99}$
GConvLSTM $4672.71_{\pm 3568.84}$ $1217.46_{\pm 929.85}$ $1206.73_{\pm 921.66}$ $695.01_{\pm 530.83}$ $527.58_{\pm 402.95}$ $326.17_{\pm 249.12}$ $505.82_{\pm 386.33}$
DyGrEncoder $3526.47_{\pm 2693.38}$ $1093.03_{\pm 834.81}$ $1365.47_{\pm 1042.9}$ $1110.66_{\pm 848.28}$ $525.05_{\pm 401.02}$ $279.39_{\pm 213.39}$ $465.82_{\pm 355.78}$
Table S11: Performance comparisons on low-frequency skill demand series with GNN-based methods.
Model Market Region L1-O L2-O R&L1-O R&L2-O Company
MAE EvolveGCNH $35.53_{\pm 27.13}$ $2.28_{\pm 1.74}$ $1.47_{\pm 1.13}$ $0.55_{\pm 0.42}$ $0.31_{\pm 0.23}$ $0.14_{\pm 0.11}$ $0.19_{\pm 0.15}$
EvolveGCNO $27.87_{\pm 21.29}$ $3.35_{\pm 2.56}$ $1.69_{\pm 1.29}$ $0.57_{\pm 0.43}$ $0.31_{\pm 0.24}$ $0.27_{\pm 0.21}$ $0.17_{\pm 0.13}$
GConvGRU $63.05_{\pm 48.16}$ $3.06_{\pm 2.34}$ $0.25_{\pm 0.19}$ $0.18_{\pm 0.14}$ $0.11_{\pm 0.09}$ $0.05_{\pm 0.04}$ $0.11_{\pm 0.09}$
TGCN $7.37_{\pm 5.63}$ $2.16_{\pm 1.65}$ $0.13_{\pm 0.1}$ $0.61_{\pm 0.46}$ $0.26_{\pm 0.2}$ $0.15_{\pm 0.12}$ $0.15_{\pm 0.11}$
GCLSTM $0.62_{\pm 0.47}$ $0.41_{\pm 0.32}$ $1.37_{\pm 1.04}$ $0.15_{\pm 0.11}$ $0.12_{\pm 0.09}$ $0.09_{\pm 0.07}$ $0.12_{\pm 0.09}$
GConvLSTM $0.79_{\pm 0.6}$ $0.43_{\pm 0.33}$ $0.2_{\pm 0.15}$ $0.13_{\pm 0.1}$ $0.1_{\pm 0.08}$ $0.06_{\pm 0.05}$ $0.13_{\pm 0.1}$
DyGrEncoder $1.04_{\pm 0.79}$ $0.47_{\pm 0.36}$ $0.24_{\pm 0.18}$ $0.13_{\pm 0.1}$ $0.15_{\pm 0.11}$ $0.09_{\pm 0.07}$ $0.14_{\pm 0.11}$
RMSE EvolveGCNH $144.53_{\pm 110.39}$ $11.08_{\pm 8.46}$ $11.65_{\pm 8.9}$ $5.52_{\pm 4.22}$ $3.48_{\pm 2.66}$ $1.7_{\pm 1.3}$ $1.81_{\pm 1.38}$
EvolveGCNO $83.5_{\pm 63.77}$ $10.36_{\pm 7.91}$ $12.82_{\pm 9.79}$ $5.74_{\pm 4.38}$ $3.8_{\pm 2.9}$ $1.95_{\pm 1.49}$ $1.65_{\pm 1.26}$
GConvGRU $63.06_{\pm 48.16}$ $3.65_{\pm 2.79}$ $1.19_{\pm 0.91}$ $0.87_{\pm 0.66}$ $1.59_{\pm 1.22}$ $0.81_{\pm 0.62}$ $1.72_{\pm 1.31}$
TGCN $21.13_{\pm 16.14}$ $8.89_{\pm 6.79}$ $1.13_{\pm 0.86}$ $3.12_{\pm 2.38}$ $2.32_{\pm 1.77}$ $1.11_{\pm 0.85}$ $1.78_{\pm 1.36}$
GCLSTM $1.63_{\pm 1.24}$ $2.48_{\pm 1.89}$ $2.8_{\pm 2.14}$ $0.86_{\pm 0.66}$ $1.61_{\pm 1.23}$ $0.83_{\pm 0.63}$ $1.73_{\pm 1.32}$
GConvLSTM $1.23_{\pm 0.94}$ $2.1_{\pm 1.6}$ $1.14_{\pm 0.87}$ $0.85_{\pm 0.65}$ $1.6_{\pm 1.22}$ $0.81_{\pm 0.62}$ $1.74_{\pm 1.33}$
DyGrEncoder $1.49_{\pm 1.14}$ $2.12_{\pm 1.62}$ $1.15_{\pm 0.88}$ $0.86_{\pm 0.66}$ $1.59_{\pm 1.22}$ $0.81_{\pm 0.62}$ $1.73_{\pm 1.32}$
SMAPE(%) EvolveGCNH $47.39_{\pm 36.19}$ $42.25_{\pm 32.27}$ $34.5_{\pm 26.35}$ $25.49_{\pm 19.47}$ $20.27_{\pm 15.48}$ $12.66_{\pm 9.67}$ $21.3_{\pm 16.27}$
EvolveGCNO $115.04_{\pm 87.86}$ $89.67_{\pm 68.48}$ $48.24_{\pm 36.84}$ $28.83_{\pm 22.02}$ $20.7_{\pm 15.81}$ $29.63_{\pm 22.63}$ $19.09_{\pm 14.58}$
GConvGRU $131.13_{\pm 100.15}$ $105.97_{\pm 80.93}$ $26.95_{\pm 20.58}$ $23.04_{\pm 17.6}$ $13.04_{\pm 9.96}$ $6.02_{\pm 4.6}$ $11.92_{\pm 9.1}$
TGCN $83.42_{\pm 63.71}$ $48.87_{\pm 37.32}$ $11.74_{\pm 8.97}$ $28.63_{\pm 21.87}$ $14.45_{\pm 11.04}$ $16.13_{\pm 12.32}$ $15.27_{\pm 11.67}$
GCLSTM $42.8_{\pm 32.69}$ $36.75_{\pm 28.07}$ $43.27_{\pm 33.05}$ $17.57_{\pm 13.42}$ $14.03_{\pm 10.72}$ $12.17_{\pm 9.3}$ $13.44_{\pm 10.26}$
GConvLSTM $53.85_{\pm 41.13}$ $36.93_{\pm 28.2}$ $21.67_{\pm 16.55}$ $14.71_{\pm 11.24}$ $10.07_{\pm 7.69}$ $7.15_{\pm 5.46}$ $14.91_{\pm 11.39}$
DyGrEncoder $62.53_{\pm 47.76}$ $39.84_{\pm 30.43}$ $26.26_{\pm 20.06}$ $15.61_{\pm 11.92}$ $17.33_{\pm 13.23}$ $12.71_{\pm 9.7}$ $16.23_{\pm 12.39}$
RRMSE(%) EvolveGCNH $66.67_{\pm 50.92}$ $67.56_{\pm 51.6}$ $66.87_{\pm 51.08}$ $67.29_{\pm 51.39}$ $73.11_{\pm 55.84}$ $73.29_{\pm 55.97}$ $138.14_{\pm 105.51}$
EvolveGCNO $66.67_{\pm 50.92}$ $67.59_{\pm 51.63}$ $66.81_{\pm 51.03}$ $67.22_{\pm 51.34}$ $69.95_{\pm 53.43}$ $71.57_{\pm 54.66}$ $119.39_{\pm 91.18}$
GConvGRU $66.46_{\pm 50.76}$ $75.73_{\pm 57.84}$ $122.73_{\pm 93.74}$ $153.14_{\pm 116.96}$ $297.91_{\pm 227.54}$ $242.77_{\pm 185.42}$ $241.43_{\pm 184.39}$
TGCN $66.51_{\pm 50.8}$ $68.02_{\pm 51.95}$ $676.51_{\pm 516.7}$ $68.73_{\pm 52.5}$ $89.88_{\pm 68.65}$ $92.87_{\pm 70.93}$ $190.19_{\pm 145.26}$
GCLSTM $60.79_{\pm 46.43}$ $103.27_{\pm 78.88}$ $70.15_{\pm 53.58}$ $169.24_{\pm 129.26}$ $359.25_{\pm 274.38}$ $214.34_{\pm 163.7}$ $272.3_{\pm 207.97}$
GConvLSTM $65.93_{\pm 50.36}$ $189.03_{\pm 144.38}$ $156.39_{\pm 119.44}$ $166.51_{\pm 127.18}$ $345.76_{\pm 264.08}$ $287.94_{\pm 219.92}$ $285.4_{\pm 217.98}$
DyGrEncoder $60.35_{\pm 46.09}$ $171.99_{\pm 131.36}$ $167.79_{\pm 128.15}$ $171.6_{\pm 131.06}$ $325.47_{\pm 248.58}$ $279.07_{\pm 213.14}$ $269.12_{\pm 205.54}$
B.4 Job Skill Demand Proportion Forecasting
In the main text, we discussed the issue of skill demand prediction. However, consider a scenario where the number of skill postings for a particular occupation is very low, leading to a low demand for that occupational skills. Nevertheless, these skills might constitute a significant portion of the profession’s core competencies. Therefore, using skill demand alone may not adequately measure the importance of these skills within the occupation. To address this, we introduce an extended dataset that includes the skill demand propotion. We define the skill demand proportion as:
$R_{s,t}^{i}=[R_{s,t,a^{i}}^{i}]_{a^{i}\in\mathcal{A}^{i}},\ \ R_{s,t,a^{i}}^{i}=\frac{\sum_{p\in\mathcal{P}_{t}}\mathbf{1}(s\in p)\cdot\mathbf{1}(a^{i}\in p)}{\sum_{p\in\mathcal{P}_{t}}\mathbf{1}(a^{i}\in p)},$ (5)
where $a^{i}\in p$ represents a job advertisement $p$ containing the attribute $a^{i}$ under granularity $i$. Similarly, we can further define skill demand proportions $R_{s,t}^{i,j,...,k}$ across multiple granularities $\{i,j,...,k\}$ by calculating:
$R_{s,t,\overline{a}}^{i,j,...,k}=\frac{\sum_{p\in\mathcal{P}_{t}}\mathbf{1}(s\in p)\cdot\mathbf{1}(a^{i}\in p\land a^{j}\in p\land...\land a^{k}\in p)}{\sum_{p\in\mathcal{P}_{t}}\mathbf{1}(a^{i}\in p\land a^{j}\in p\land...\land a^{k}\in p)},$ (6)
where $\overline{a}=\{a^{i},a^{j},...,a^{k}\}$, $a^{i}\in\mathcal{A}^{i},a^{j}\in\mathcal{A}^{j},...,a^{k}\in\mathcal{A}^{k}$, and $R^{i,j,\ldots,k}_{s,t}\in\mathbb{R}^{|\mathcal{A}^{i}||\mathcal{A}^{j}|\ldots|\mathcal{A}^{k}|}$.
Results
We continue to utilize the benchmark models described in the main text for this task, and the results, as shown in Table S12, lead to the following conclusions: Firstly, the best-performing model on the task of forecasting the proportion of skill demand is Koopa. This model, integrating time series decomposition and Fourier transformations, effectively captures the distribution changes in demand proportions. Secondly, there is a significant variation in performance across models in this task. For example, models like DLinear perform poorly on this task, though they are reasonably effective in skill demand forecasting. We analyze that predicting percentages is distinct from forecasting skill demand, as percentage predictions are also influenced by the demand for other skills at the same granularity. Therefore, simple linear models are not advantageous for capturing the complex interrelations and influences among multiple pieces of information.
Table S12: Performance comparisons on skill demand proportion forecasting.
Model Market Region L1-O L2-O R&L1-O R&L2-O Company
MAE(%) LSTM $0.14\pm_{0.0}$ $0.49\pm_{0.01}$ $1.21\pm_{0.03}$ $2.26\pm_{0.05}$ $2.43\pm_{0.03}$ $3.51\pm_{0.06}$ $2.69\pm_{0.03}$
SegRNN $0.15\pm_{0.02}$ $1.25\pm_{0.02}$ $2.63\pm_{0.11}$ $4.2\pm_{0.04}$ $6.6\pm_{1.12}$ $10.16\pm_{2.44}$ $5.52\pm_{0.53}$
CHGH $0.1\pm_{0.01}$ $0.19\pm_{0.0}$ $0.35\pm_{0.0}$ $0.64\pm_{0.01}$ $0.7\pm_{0.0}$ $0.91\pm_{0.07}$ $1.03\pm_{0.01}$
PreDyGAE $0.08\pm_{0.01}$ $0.09\pm_{0.0}$ $0.12\pm_{0.02}$ $0.23\pm_{0.02}$ $0.18\pm_{0.0}$ $0.27\pm_{0.15}$ $0.29\pm_{0.16}$
Transformer $0.62\pm_{0.02}$ $3.89\pm_{0.02}$ $10.87\pm_{0.05}$ $20.58\pm_{0.01}$ $22.02\pm_{0.01}$ $31.26\pm_{0.02}$ $23.47\pm_{0.01}$
Autoformer $1.29\pm_{0.14}$ $8.84\pm_{0.15}$ $24.5\pm_{0.22}$ $43.99\pm_{0.94}$ $47.53\pm_{0.67}$ $66.32\pm_{1.02}$ $50.02\pm_{0.66}$
Informer $0.55\pm_{0.02}$ $3.83\pm_{0.03}$ $10.86\pm_{0.03}$ $20.6\pm_{0.01}$ $22.04\pm_{0.01}$ $31.23\pm_{0.01}$ $23.48\pm_{0.0}$
Reformer $0.56\pm_{0.01}$ $3.85\pm_{0.01}$ $10.9\pm_{0.03}$ $20.65\pm_{0.02}$ $22.01\pm_{0.01}$ $31.25\pm_{0.01}$ $23.48\pm_{0.0}$
FEDformer $1.42\pm_{0.04}$ $8.58\pm_{0.18}$ $25.11\pm_{0.59}$ $44.15\pm_{0.6}$ $47.07\pm_{0.66}$ $65.95\pm_{0.23}$ $50.0\pm_{0.07}$
NStransformer $0.06\pm_{0.0}$ $0.09\pm_{0.0}$ $0.11\pm_{0.0}$ $0.19\pm_{0.0}$ $0.24\pm_{0.0}$ $0.34\pm_{0.0}$ $0.34\pm_{0.0}$
PatchTST $0.06\pm_{0.0}$ $0.09\pm_{0.01}$ $0.11\pm_{0.02}$ $0.18\pm_{0.04}$ $0.23\pm_{0.05}$ $0.33\pm_{0.07}$ $0.33\pm_{0.06}$
DLinear $0.27\pm_{0.15}$ $1.65\pm_{1.07}$ $4.63\pm_{3.09}$ $8.76\pm_{5.86}$ $9.4\pm_{6.26}$ $13.33\pm_{8.89}$ $10.05\pm_{6.65}$
TSMixer $0.85\pm_{0.03}$ $3.15\pm_{0.69}$ $7.57\pm_{1.43}$ $18.88\pm_{0.49}$ $15.12\pm_{3.85}$ $18.76\pm_{2.85}$ $17.77\pm_{1.58}$
FreTS $0.12\pm_{0.0}$ $0.31\pm_{0.04}$ $0.69\pm_{0.11}$ $1.26\pm_{0.21}$ $1.37\pm_{0.22}$ $1.94\pm_{0.32}$ $1.54\pm_{0.24}$
FiLM $0.06\pm_{0.01}$ $0.09\pm_{0.01}$ $0.12\pm_{0.01}$ $0.19\pm_{0.02}$ $0.25\pm_{0.03}$ $0.35\pm_{0.05}$ $0.34\pm_{0.05}$
Koopa $0.06\pm_{0.0}$ $0.08\pm_{0.0}$ $0.09\pm_{0.0}$ $0.15\pm_{0.01}$ $0.2\pm_{0.01}$ $0.27\pm_{0.02}$ $0.28\pm_{0.01}$
RMSE(%) LSTM $0.61\pm_{0.01}$ $1.74\pm_{0.04}$ $2.89\pm_{0.08}$ $4.06\pm_{0.09}$ $4.16\pm_{0.05}$ $5.21\pm_{0.1}$ $4.43\pm_{0.05}$
SegRNN $1.32\pm_{0.17}$ $4.98\pm_{0.11}$ $6.49\pm_{0.22}$ $7.7\pm_{0.09}$ $11.23\pm_{1.78}$ $14.58\pm_{3.23}$ $9.18\pm_{0.83}$
CHGH $0.36\pm_{0.01}$ $0.49\pm_{0.0}$ $0.7\pm_{0.0}$ $1.12\pm_{0.02}$ $1.14\pm_{0.0}$ $2.11\pm_{0.32}$ $2.31\pm_{0.07}$
PreDyGAE $0.23\pm_{0.0}$ $0.29\pm_{0.0}$ $0.36\pm_{0.0}$ $0.85\pm_{0.0}$ $0.91\pm_{0.0}$ $1.15\pm_{1.05}$ $1.29\pm_{0.72}$
Transformer $6.22\pm_{0.12}$ $16.65\pm_{0.08}$ $27.9\pm_{0.14}$ $38.45\pm_{0.03}$ $39.77\pm_{0.01}$ $47.4\pm_{0.02}$ $40.97\pm_{0.01}$
Autoformer $15.11\pm_{2.51}$ $38.94\pm_{0.61}$ $63.72\pm_{0.49}$ $82.56\pm_{1.83}$ $86.21\pm_{1.22}$ $100.99\pm_{1.58}$ $87.83\pm_{1.13}$
Informer $5.63\pm_{0.15}$ $16.41\pm_{0.13}$ $27.89\pm_{0.05}$ $38.5\pm_{0.03}$ $39.79\pm_{0.01}$ $47.36\pm_{0.01}$ $40.98\pm_{0.0}$
Reformer $5.62\pm_{0.05}$ $16.45\pm_{0.05}$ $27.96\pm_{0.11}$ $38.58\pm_{0.02}$ $39.76\pm_{0.03}$ $47.38\pm_{0.02}$ $40.98\pm_{0.02}$
FEDformer $17.24\pm_{0.68}$ $37.52\pm_{0.77}$ $64.93\pm_{1.42}$ $82.85\pm_{1.16}$ $85.38\pm_{1.21}$ $100.35\pm_{0.34}$ $87.74\pm_{0.1}$
NStransformer $0.21\pm_{0.0}$ $0.28\pm_{0.0}$ $0.35\pm_{0.0}$ $0.77\pm_{0.01}$ $0.86\pm_{0.0}$ $1.69\pm_{0.0}$ $1.24\pm_{0.0}$
PatchTST $0.21\pm_{0.01}$ $0.28\pm_{0.02}$ $0.35\pm_{0.03}$ $0.75\pm_{0.04}$ $0.84\pm_{0.06}$ $1.65\pm_{0.1}$ $1.22\pm_{0.07}$
DLinear $2.21\pm_{1.49}$ $6.05\pm_{4.1}$ $10.27\pm_{6.97}$ $14.16\pm_{9.59}$ $14.64\pm_{9.91}$ $17.51\pm_{11.77}$ $15.11\pm_{10.19}$
TSMixer $10.5\pm_{0.04}$ $14.32\pm_{2.43}$ $21.09\pm_{3.64}$ $34.92\pm_{1.37}$ $25.27\pm_{5.66}$ $26.48\pm_{3.26}$ $31.07\pm_{2.82}$
FreTS $0.47\pm_{0.02}$ $0.96\pm_{0.07}$ $1.56\pm_{0.12}$ $2.22\pm_{0.17}$ $2.29\pm_{0.17}$ $2.99\pm_{0.2}$ $2.54\pm_{0.17}$
FiLM $0.21\pm_{0.0}$ $0.28\pm_{0.0}$ $0.35\pm_{0.02}$ $0.76\pm_{0.04}$ $0.85\pm_{0.06}$ $1.67\pm_{0.11}$ $1.22\pm_{0.05}$
Koopa $0.21\pm_{0.0}$ $0.27\pm_{0.0}$ $0.33\pm_{0.01}$ $0.72\pm_{0.01}$ $0.79\pm_{0.02}$ $1.57\pm_{0.04}$ $1.17\pm_{0.01}$
SMAPE(%) LSTM $0.26\pm_{0.0}$ $0.92\pm_{0.02}$ $2.25\pm_{0.05}$ $4.21\pm_{0.09}$ $4.52\pm_{0.05}$ $6.5\pm_{0.11}$ $5.0\pm_{0.05}$
SegRNN $0.27\pm_{0.03}$ $2.09\pm_{0.03}$ $4.55\pm_{0.17}$ $7.38\pm_{0.06}$ $11.1\pm_{1.63}$ $16.79\pm_{3.54}$ $9.58\pm_{0.83}$
CHGH $0.19\pm_{0.01}$ $0.36\pm_{0.0}$ $0.69\pm_{0.0}$ $1.24\pm_{0.02}$ $1.36\pm_{0.0}$ $1.78\pm_{0.01}$ $1.77\pm_{0.06}$
PreDyGAE $0.09\pm_{0.0}$ $0.19\pm_{0.0}$ $0.22\pm_{0.0}$ $0.28\pm_{0.0}$ $0.39\pm_{0.0}$ $0.64\pm_{0.27}$ $0.57\pm_{0.3}$
Transformer $0.81\pm_{0.02}$ $4.78\pm_{0.01}$ $13.28\pm_{0.04}$ $25.1\pm_{0.0}$ $26.89\pm_{0.01}$ $38.12\pm_{0.01}$ $28.71\pm_{0.01}$
Autoformer $1.13\pm_{0.02}$ $7.42\pm_{0.05}$ $20.74\pm_{0.1}$ $38.35\pm_{0.38}$ $41.28\pm_{0.29}$ $58.08\pm_{0.43}$ $43.84\pm_{0.29}$
Informer $0.75\pm_{0.02}$ $4.74\pm_{0.02}$ $13.27\pm_{0.03}$ $25.12\pm_{0.0}$ $26.9\pm_{0.01}$ $38.1\pm_{0.01}$ $28.72\pm_{0.0}$
Reformer $0.75\pm_{0.01}$ $4.75\pm_{0.01}$ $13.31\pm_{0.02}$ $25.16\pm_{0.02}$ $26.87\pm_{0.01}$ $38.12\pm_{0.01}$ $28.72\pm_{0.0}$
FEDformer $1.17\pm_{0.01}$ $7.35\pm_{0.08}$ $21.02\pm_{0.24}$ $38.42\pm_{0.24}$ $41.09\pm_{0.28}$ $57.94\pm_{0.1}$ $43.84\pm_{0.03}$
NStransformer $0.11\pm_{0.0}$ $0.17\pm_{0.0}$ $0.22\pm_{0.0}$ $0.35\pm_{0.0}$ $0.45\pm_{0.0}$ $0.61\pm_{0.0}$ $0.63\pm_{0.0}$
PatchTST $0.11\pm_{0.01}$ $0.17\pm_{0.02}$ $0.21\pm_{0.05}$ $0.34\pm_{0.09}$ $0.44\pm_{0.1}$ $0.59\pm_{0.14}$ $0.61\pm_{0.11}$
DLinear $0.43\pm_{0.19}$ $2.58\pm_{1.39}$ $7.19\pm_{4.0}$ $13.59\pm_{7.59}$ $14.59\pm_{8.11}$ $20.67\pm_{11.52}$ $15.62\pm_{8.62}$
TSMixer $0.92\pm_{0.03}$ $4.05\pm_{0.67}$ $10.03\pm_{1.36}$ $23.97\pm_{1.21}$ $21.46\pm_{4.26}$ $27.6\pm_{3.5}$ $23.83\pm_{1.45}$
FreTS $0.22\pm_{0.01}$ $0.59\pm_{0.07}$ $1.31\pm_{0.21}$ $2.4\pm_{0.41}$ $2.63\pm_{0.43}$ $3.69\pm_{0.62}$ $2.95\pm_{0.46}$
FiLM $0.11\pm_{0.02}$ $0.17\pm_{0.02}$ $0.22\pm_{0.02}$ $0.36\pm_{0.03}$ $0.46\pm_{0.06}$ $0.63\pm_{0.09}$ $0.64\pm_{0.09}$
Koopa $0.1\pm_{0.0}$ $0.15\pm_{0.01}$ $0.18\pm_{0.01}$ $0.28\pm_{0.01}$ $0.37\pm_{0.02}$ $0.49\pm_{0.03}$ $0.52\pm_{0.03}$
RRMSE(%) LSTM $43.76\pm_{0.35}$ $81.48\pm_{0.6}$ $85.53\pm_{0.64}$ $91.34\pm_{0.35}$ $93.23\pm_{0.15}$ $97.48\pm_{0.09}$ $96.6\pm_{0.07}$
SegRNN $72.41\pm_{4.5}$ $96.92\pm_{0.17}$ $96.4\pm_{0.21}$ $96.99\pm_{0.11}$ $98.79\pm_{0.46}$ $99.24\pm_{0.25}$ $98.15\pm_{0.28}$
CHGH $27.4\pm_{0.7}$ $37.66\pm_{0.07}$ $37.45\pm_{0.03}$ $53.25\pm_{0.58}$ $58.32\pm_{0.03}$ $80.91\pm_{5.07}$ $70.89\pm_{2.0}$
PreDyGAE $20.64\pm_{0.0}$ $28.83\pm_{0.0}$ $40.02\pm_{0.01}$ $40.12\pm_{0.0}$ $47.5\pm_{0.0}$ $76.9\pm_{42.81}$ $66.68\pm_{33.49}$
Transformer $97.99\pm_{0.07}$ $99.73\pm_{0.01}$ $99.8\pm_{0.0}$ $99.89\pm_{0.0}$ $99.92\pm_{0.0}$ $99.97\pm_{0.0}$ $99.96\pm_{0.0}$
Autoformer $99.63\pm_{0.12}$ $99.95\pm_{0.0}$ $99.96\pm_{0.0}$ $99.97\pm_{0.0}$ $99.98\pm_{0.0}$ $99.99\pm_{0.0}$ $99.98\pm_{0.0}$
Informer $97.59\pm_{0.12}$ $99.72\pm_{0.0}$ $99.8\pm_{0.0}$ $99.89\pm_{0.0}$ $99.92\pm_{0.0}$ $99.97\pm_{0.0}$ $99.96\pm_{0.0}$
Reformer $97.57\pm_{0.03}$ $99.72\pm_{0.01}$ $99.81\pm_{0.0}$ $99.89\pm_{0.0}$ $99.92\pm_{0.0}$ $99.97\pm_{0.0}$ $99.96\pm_{0.0}$
FEDformer $99.75\pm_{0.02}$ $99.95\pm_{0.0}$ $99.96\pm_{0.0}$ $99.97\pm_{0.0}$ $99.98\pm_{0.0}$ $99.99\pm_{0.0}$ $99.98\pm_{0.0}$
NStransformer $16.44\pm_{0.44}$ $22.17\pm_{0.31}$ $19.66\pm_{0.16}$ $37.34\pm_{0.26}$ $45.81\pm_{0.26}$ $75.98\pm_{0.13}$ $61.84\pm_{0.13}$
PatchTST $16.32\pm_{0.88}$ $21.85\pm_{1.31}$ $19.5\pm_{1.44}$ $36.66\pm_{1.59}$ $44.97\pm_{1.95}$ $75.59\pm_{0.49}$ $61.55\pm_{1.34}$
DLinear $78.33\pm_{15.19}$ $95.41\pm_{3.54}$ $96.69\pm_{2.57}$ $97.94\pm_{1.58}$ $98.55\pm_{1.11}$ $99.37\pm_{0.44}$ $98.87\pm_{0.84}$
TSMixer $99.29\pm_{0.02}$ $99.61\pm_{0.13}$ $99.64\pm_{0.11}$ $99.86\pm_{0.02}$ $99.8\pm_{0.11}$ $99.85\pm_{0.0}$ $99.91\pm_{0.05}$
FreTS $34.81\pm_{1.43}$ $61.32\pm_{3.03}$ $66.35\pm_{3.08}$ $77.25\pm_{2.53}$ $81.41\pm_{2.37}$ $92.32\pm_{1.5}$ $90.05\pm_{1.64}$
FiLM $16.49\pm_{0.39}$ $21.81\pm_{0.76}$ $19.48\pm_{0.05}$ $36.44\pm_{0.97}$ $44.79\pm_{2.15}$ $75.91\pm_{10.46}$ $61.27\pm_{8.05}$
Koopa $16.47\pm_{0.04}$ $21.47\pm_{0.18}$ $18.71\pm_{0.07}$ $35.58\pm_{0.07}$ $43.65\pm_{0.38}$ $75.85\pm_{2.58}$ $61.26\pm_{1.98}$
Appendix C Data Structure and Components
Our dataset comprises five components for each granularity level: job skill demand sequences, job skill demand proportion sequences, ID mapping index, the indexes of skills with structural breaks, and skill co-occurrence graph. Each component is structured as follows: (1) Job Skill Demand Sequences: These are presented in tabular files, where each row represents a specific skill, and each column corresponds to a different time slice (month). Each cell within the table contains a numerical value that reflects the demand for the respective skill during that month. (ii) Job Skill Demand Proportion Sequences: This component is also formatted in tabular files similar to the skill demand sequences. However, each cell in these tables displays a value between 0 and 1, representing the proportion of demand defined in Eq 6. This provides a normalized view of skill demand across different granularities. (iii) ID mapping index: In the dataset, various elements such as regions, occupations, companies, and skills are represented using unique identifiers (IDs) for the convenience of experimentation and analysis. An index table is provided that maps each ID to the actual names of regions, occupations, and skills, facilitating clear and effective data interpretation. The names of companies, however, are withheld due to potential privacy concerns. This selective anonymization ensures that while company-related data remains confidential, researchers can still access necessary details about other variables. During the review period, we open part of id-to-entity mappings to illustrate its use and structure. (iv) Indexes of skills with structural breaks: In the provided dataset, data concerning skills that have experienced structural breaks are organized in JSON format. Each granularity level is represented by a separate JSON file, which contains a list of indexes. These indexes correspond to the skills that have undergone structural breaks and can be directly mapped to the skill indexes in the skill demand sequences. The purpose of supplying this data is to facilitate research on the demand trends of skills that have exhibited structural breaks, enabling a detailed analysis of their demand dynamics over time. (v) Skill Co-occurrence Graph: This data is provided as a set of triples (skill ID_1, skill ID_2, frequency of co-occurrence), forming a collection that outlines the co-occurrence relationships between skills. Each triple indicates how frequently two skills are mentioned or required together within the job advertisements in the training data, serving as a prior knowledge graph to enhance predictive modeling by capturing relationships between skills.